### v1

# Atlas: A Framework for ML Lifecycle Provenance & Transparency

Abstract
The rapid adoption of open source machine learning (ML) datasets and models exposes today’s AI applications to critical risks like data poisoning and supply chain attacks across the ML lifecycle. With growing regulatory pressure to address these issues through greater transparency, ML model vendors face challenges balancing these requirements against confidentiality for data and intellectual property needs.
We propose Atlas, a framework that enables fully attestable ML pipelines. Atlas leverages open specifications for data and software supply chain provenance to collect verifiable records of model artifact authenticity and end-to-end lineage metadata. Atlas combines trusted hardware and transparency logs to enhance metadata integrity, preserve data confidentiality, and limit unauthorized access during ML pipeline operations, from training through deployment.
Our prototype implementation of Atlas integrates several open-source tools to build an ML lifecycle transparency system, and assess the practicality of Atlas through two case study ML pipelines.
1 Introduction
In recent years, machine learning (ML) models, have become increasingly popular. The pervasive use of large language models (LLMs), in particular, and multi-stakeholder involvement in model creation and deployment exacerbate security and privacy risks. These considerations are emphasized by the global nature and the complexity of large-scale ML deployments with different lifecycle stages:
  1. 1)
    Collection and sanitation of a training dataset from several public and proprietary sources.
  2. 2)
    Provisioning of the training environment (hardware and software).
  3. 3)
    Execution of training across many data centers.
  4. 4)
    Construction of a testing dataset from several sources, and the evaluation.
  5. 5)
    Deployment and use of the model for inference that is compliant with local laws or corporate policies.
Each of these stages is vulnerable to malicious or dishonest parties. For example, data can be poisoned [1, 2] during collection or training. Service providers executing outsourced training can shorten or omit critical steps to reduce their cost. Model providers can serve smaller models in SaaS, or even distribute malicious ones.
On the other hand, responsible model builders and other stakeholders may be incentivised or required to provide security and trust guarantees. They may want to prove low bias in their training data, offer easily verifiable performance claims, or guarantee end-to-end integrity of the model creation in high risk domains.
To address these challenges, it is necessary to guarantee the integrity of the entire ML lifecycle – beginning with the data, through the training, and finally, the evaluation and deployment. Was the data modified? Did the hardware and software environment adhere to the specification? Did the contractor follow the specified training procedure? Can I trust the evaluation? How can I guarantee that I am interacting with the intended model? These are example questions that showcase the breadth of the involved challenges that must be tackled to provide end-to-end security.
In this work, we introduce Atlas, a framework for enhancing the security and transparency of the lifecycle of ML models. Atlas establishes the baseline of fundamental components and capabilities needed for comprehensive provenance tracking at each stage of the ML lifecycle. Subsequently, Atlas defines the core integrity requirements for verifiable ML lifecycle transparency. We provide a reference implementation that instantiates Atlas using hardware-based security mechanisms – with trusted execution environment (TEE), including attestations.
We claim the following contributions:
  1. 1.
    We introduce Atlas, a framework designed for end-to-end ML lifecycle transparency.
  2. 2.
    We instantiate Atlas using TEEs and metadata-based provenance tracking.
  3. 3.
    We evaluate our Atlas prototype through two case studies: 1) fine-tuning of a BERT model [3, 4]; 2) fine-tuning of a bge-reranker model [5] .
2 Background & Related Work
2.1 Data Provenance & Authenticity
While not a new concern, the authenticity, provenance and attribution of media has received renewed attention in recent years due to the wide online proliferation of images, videos and text that have been manipulated or forged using generative ML models [6, 7].
One common approach for addressing digital image and video manipulation involves cryptographic techniques such as hashing, digital signatures and blockchains, which enable provenance authentication, tamper detection, and checking for divergent views of the content [8].
Most notably, the Coalition for Content Provenance and Authenticity (C2PA) specification [9, 10] introduced by the Content Authenticity Initiative (CAI) [11] and Project Origin [12] provides cryptographic binding of authenticity claims, verifiable assertions about content origin and modifications, and tamper-evident audit trails [13].
Initially launched as a response to the growing challenge of deepfakes  [11] and digital content manipulation, the C2PA standard first gained traction in digital photography and journalism workflows. The standard’s extensible and interoperable metadata format has made C2PA an attractive option for ML dataset and model provenance tracking [14].
Other prior work in this space builds upon digitally signed metadata by leveraging distributed ledger technologies to create transparent and immutable content or provenance records. Examples of this approach include AMP [7], the News Provenance Project [15] and Collomosse et al. [14].
Atlas borrows techniques from C2PA and similar approaches, but integrates them directly into the ML system to track model artifact provenance closer to where the transformations occur to achieve higher metadata fidelity, rather than applying them in an ad hoc fashion after a particular dataset or model artifact has been created.
The Open Source Security Foundation (OpenSSF) Model Signing project [16] is a parallel effort in this space focusing specifically on cryptographic signing of model artifacts to ensure their authenticity and integrity. As such, it serves as a crucial building block but requires integration with other tools for more comprehensive supply chain security. While Model Signing might integrate with Atlas, additional mechanisms would be needed to address our goal of tracking full model lineage.
LakeFS [17] combines Git-like semantics with concepts from object stores such as S3 to provide a version control system for data, including ML datsets. Thus, LakeFS aims to capture data lineage by tracking changes to stored data over time, and allowing ML applications to reference specific versions of the stored data. This approach is meant to integrate with existing first-party data processing pipelines, but does not facilitate verification of data provenance by downstream consumers. Atlas’ metadata centered approach, on the other hand, enables first- and third-party ML dataset consumers to track changes and check their provenance, even when they may not have direct access to the data.
2.2 Supply Chain Integrity
While the ML supply chain is an emerging area of research, software supply chain security and integrity has been the subject of much prior work. Thus, several of these approaches are being explored for their applicability to enhancing the integrity of the ML model lifecycle.
Bills of Materials (BOM) have been employed to document the list of components of a hardware or software product for over three decades [18]. Due to recent regulations [19, 20], Software BOM (SBOM) have been the focus of many industry and academic efforts seeking to facilitate tracking software dependencies and other metadata to detect supply chain threats [21].
The AIBOM framework [22, 23] incorporates concepts from traditional software supply chain management while focusing on intended ML model use cases. Like SBOM, AIBOM provide a mechanism for tracking model software dependencies and maintaining model metadata. At the time of writing, we are not aware of any frameworks other than Atlas that utilize any sort of BOM data format to track ML model components.
Complementing BOM, a number of efforts like OpenSSF Supply Chain Levels for Software Artifacts (SLSA) [24] and SPDX Build [25] have sought to collect build provenance, i.e., metadata describing the process through which a particular artifact was produced. With its incremental guidance on implementing integrity measures for software builds and its build provenance data format, the SLSA framework [24] is gaining traction among software vendors, and also beginning to see some use for ML model fine tuning [26]. Atlas’ configurable metadata collection capabilities would enable us to collect build provenance at any stage of the ML model lifecycle, providing a more comprehensive view.
A number of frameworks for capturing and verifying a variety of security claims and metadata about the supply chain have been proposed. in-toto [27] collects authenticated claims across supply chain steps, including SBOM and SLSA metadata. In particular, in-toto enables software development pipeline owners and downstream artifact consumers to specify end-to-end supply chain policies, and validate that only the expected parties carried out specific steps in the supply chain, and artifacts underwent transformations in the expected order. Given recent and upcoming enhancements that further generalize the framework, in-toto may be a suitable option for specifying and verifying end-to-end ML model lifecycle integrity policies in Atlas.
Sigstore [28] provides a transparency log-based infrastructure for issuing signing credentials and validating digital signatures on supply chain artifacts and metadata. Similarly, Supply Chain Integrity, Transparency and Trust (SCITT) [29] is an architecture for implementing distributed ledger to provide global visibility and auditing for supply chain operations and claims. The SCITT architecture also includes confidential computing technologies to help ensure that only authorized parties submit claims to the transparency ledger.
Atlas builds upon several of these supply chain integrity and transparency approaches, but seeks to support multiple types of artifact provenance, supply chain metadata and integrity verification mechanisms all within a customizable, integrated framework designed for the ML model lifecycle.
2.3 Model Lineage Tracking
The EQTY Lineage Explorer [30] provides granular tracking of model artifacts throughout the training process. The tool captures relationships between datasets, model checkpoints and hyperparameters, enabling developers to trace model evolution. Unlike Atlas, it lacks cryptographic guarantees for artifact authenticity and focuses primarily on manually collected development-time lineage, rather than automatically capturing and linking information across the entire ML lifecycle.
ML experiment trackers like Weights and Biases [31], Neptune [32] and Kubeflow Pipelines [33] offer detailed run-time logging of model metadata about training runs, metrics, and model artifacts. While comprehensive for collaborative ML application development, these tools do not necessarily integrate transparently with common ML frameworks, and they typically provide only unauthenticated metadata that makes it less appropriate for establishing trust in a particular ML pipeline. Atlas, on the other hand, focuses making model lineage verifiable and on enabling adoption by supporting integration into existing open ML frameworks.
2.4 Hardware-Based Security for ML
Recent developments in trusted execution environments (TEEs) technologies have made it more practical to run large-scale systems and workloads. Thus, a growing number of proposals leverage TEEs to harden a variety of ML systems through hardware-enforced isolation and integrity verification mechanisms.
Chrapek et al. [34] use TEEs in a large language model (LLM) fine-tuning system running inside secure enclaves that help protect LLM code and data while in use. To this end, they evaluate their LLM system using the process-level Intel®Software Guard eXtensions (Intel®SGX) and virtual machine-level Intel®Trust Domain eXtensions (Intel®TDX) TEEs, showing that they are able to maintain practical performance overhead of less than 10% for models like Llama2 in both TEE configurations.
Laminator [35] demonstrates the application of TEEs to ML property attestation for ML model regulatory compliance verification. Specifically, this work leverages TEEs’ hardware-assisted remote attestation and cryptographic binding capabilities to generate verifiable ML model property cards. This work is complementary to Atlas and may enable us to extend our framework to support such ML property attestations.
Mo et al.’s survey [36] evaluates 38 works that use various TEE implementations to enhance the privacy and integrity of ML training and inference operations. The survey highlights several gaps that existing TEE-based frameworks do not address, including the protection of a full ML lifecycle, which is the primary focus of Atlas.
3 System Overview & Threat Model
3.1 Terminology
In Atlas, an ML model is composed of a number of artifacts that include the training dataset, ML algorithm, ML framework (e.g., PyTorch), model configuration (e.g., hyperparameters, weights), and metadata (e.g., license).
The ML lifecycle consists of various stages, including data preparation, training, evaluation and deployment. A common synonym for ML lifecycle is “ML supply chain”, so we use these terms interchangeably in the paper.
The ML pipeline defines the sequence of operations or a workflow that transforms or operates on a model artifact at a particular stage of the ML lifecycle [37]. To support the standardization and repeatability of a specific workflow, the pipeline is also used to facilitate workflow management and automation, e.g., training or evaluation pipelines.
The ML system is the set of hardware and software components that implement and execute an ML pipeline. For example, an ML system for training may include orchestration tools, an authentication service, storage systems, automation infrastructure, and specialized compute hardware (e.g., GPUs, TPUs, or custom accelerators).
3.2 System Model
In Atlas, we target the multi-stakeholder ML model lifecycle setting.
3.2.1 Stakeholders
The Artifact Producers are individuals and organizations that create ML model artifacts to provide or sell them to other parties. Since the artifacts may represent intellectual property and/or make use of personally identifying information (PII), producers have business and regulatory reasons to preserve the confidentiality of their artifacts. To save costs, artifact producers often outsource the operation of ML systems to external MLaaS providers.
ML-as-a-Service (MLaaS) providers operate and maintain the compute infrastructure needed to run ML systems on behalf of artifact producers. MLaaS providers may offer ML-specific services that leverage general-purpose systems (e.g., [38, 39, 40]), or provide special-purpose systems like orchestrators [33] or CI/CD (e.g. [41]) that can be used to build and run third-party ML systems.
A Hub is a system that stores and distributes model artifacts. Thus, model pipelines typically ingest and output artifacts to and from hubs during their execution, enabled by interfaces exposed by MLaaS providers.
Hubs may be operated by artifact producers themselves or by third-party vendors, containing open or closed source artifacts (e.g.,HuggingFace [42], PyTorch Hub [43]).
A Transparency Service in Atlas is responsible for generating, storing and distributing the information necessary to verify the authenticity and integrity of model artifacts, ML systems and lifecycles as a whole. We envision model vendors and independent parties to operate transparency services in practice.
Verifiers are entities that seek to evaluate or audit a particular ML model’s lifecycle with the goal of checking whether the model it produced was tampered with in an unintended or malicious way. To this end, verifiers in Atlas define a set of expectations about a producer’s model artifacts, MLaaS providers, and their ML system components. In practice, model users, vendors or regulatory entities may act as verifiers.
A Model User interacts with a model in an inferencing pipeline, or in a downstream ML pipeline as a dependency, such as a fine-tuning or evaluation pipeline (see §3.2.2).
3.2.2 ML Lifecycle
In Atlas, we consider four common high-level stages in the ML lifecycle. Each builds upon the outputs and feedback from the others, forming a continuous cycle that allows models to evolve and adapt based on real-world usage.
1. Data processing: Raw data is collected, curated, processed into smaller units (e.g., tokens) and collated into a structure ingestable during training.
2. Training: a training algorithm processes a given dataset using an ML framework. The output of the training pipeline is an ML model that may undergo further fine tuning and evaluation, prior to being deployed for inferencing.
3. Evaluation: following training, model properties like its performance and accuracy are checked and further adjusted, prior to production release and deployment.
4. Deployment: Once an ML model has been trained and evaluated, it is deployed to a production system configured for inferencing. The new data obtained from clients during inference is sent back to a data processing pipeline to enhance the training dataset, and the model.
3.3 Threat Model
Atlas is concerned with addressing risks to ML model artifacts introduced via the ML lifecycle, i.e., attack vectors in third-party MLaaS and hubs across different ML pipelines that lead to compromised ML artifacts. Specifically, while transparency services, verifiers and model users are trusted in Atlas, we consider threats by MLaaS providers, hubs and artifact producers.
The adversary’s goal is to produce a tampered artifact or model, e.g., containing a hidden malicious component, so that a trusted transparency service generates a legitimate signature on the artifact or its metadata.
Compromised MLaaS providers and hubs may involve malicious insiders, or external adversaries seeking to subvert these systems by exploiting vulnerable components. Given their central position, MLaaS providers and hubs may thus be able to compromise model integrity at various stages of the ML lifecycle.
For example, a malicous MLaaS provider can poison the training data during the curation step of the data processing stage leading to backdoors [2, 1]. A malicious hub may, for instance, present a dataset or model with a mismatched signature (e.g., to a different model, or any of its component artifacts) to a model user or MLaaS provider, so that pipelines in subsequent stages of the ML lifecycle may ingest compromised dependencies.
As a result, these compromises may propagate through the ML lifecycle if they go undetected, ultimately leading to vulnerable ML models at the deployment stage. This risk is exacerbated if a hub colludes with an MLaaS providers to introduce or accept compromised ML pipeline inputs.
Artifact producers, on the other hand, may seek to compromise ML models in order to bypass regulatory requirements, introduce exploitable vulnerabilities or steal private information for profit (e.g, [44]). Thus, producers may collude with other untrusted participants, or inaccurately declare their dependencies leading to intentional omissions or misrepresentations their artifacts, to surreptitiously undermine the integrity of their artifacts.
3.3.1 Trusted Parties
As mentioned, Atlas considers the transparency services, verifiers and model users in an ML lifecycle to be trusted. We make the following assumptions about the security of the systems underlying and enabling the participants: 1) the hardware and cryptographic primitives are implemented correctly and do not contain known vulnerabilities; 2) a separate PKI system and organizations representing the participants follow best practices for key management, network security and access control.
This allows us to trust that transparency services correctly implement Atlas attestation mechanisms, detect tampering with them by a malicious MLaaS provider and maintain accurate golden values of model artifacts. Verifiers are trusted to properly evaluate Atlas attestations, including their digital signatures, against pre-specified expectations, and to publish a record of their evaluations.
3.3.2 Out of Scope
We consider the following threats to be out of scope for Atlas. While a critical threat to deployed AI applications, inference time black-box attacks (e.g., evasion attacks, model extraction, membership inference) are caused by malicious users, which we do not address; solutions proposed to reduce this particular risk [45, 46, 47] are complementary to our work.
Side-channel attacks against hardware enclaves, physical attacks on hardware infrastructure, and network-level denial of service attacks are also beyond the scope of Atlas. These attacks are the subject of a large body of prior work [36], and we acknowledge that these complementary security measures could be added to deployments of Atlas.
3.4 Design Requirements
We define the following integrity and operational requirements for Atlas:
R1: Artifact tampering is detectable. To provide model artifact authenticity, Atlas must enable verifiers to detect unexpected modifications to model artifacts by malicious MLaaS providers or hubs.
R2: Every model transformation is attested. Because adversaries may seek to tamper with model artifacts after they are produced by a pipeline, every model transformation in Atlas must produce an authenticated record as evidence for the process.
R3: Verifiable model lineage. Atlas verifiers must be able to detect unauthorized operators and unintended or malicious changes to the expected stages of the lifecycle (e.g., pipelines operating out of order, or being omitted).
R4: Strongly isolated ML systems. To reduce the risks of tampering with a pipeline during its execution, Atlas must restrict access to its ML system by unauthorized entities, and contain compromises from propagating beyond the execution environment.
R5: Controlled artifact disclosure. Atlas must address data and model producer concerns about PII and intellectual property confidentiality by enabling participants to control which parties and systems are able to access datasets and model components.
R6: Pipeline agnostic. To facilitate adoption, Atlas must be agnostic to any ML pipeline that integrates it.
R7: Efficiency. We seek to minimize the computational and storage overheads incurred by Atlas to enable the implementation and deployment of Atlas in ML systems using commodity platforms and services.
4 Atlas Framework
Atlas introduces two core components to the ML lifecycle: 1) the transparency service interacting with MLaaS providers; 2) the verification service that model users and verifiers use to verify integrity and provenance. Fig. 1 depicts an example ML model lifecycle with Atlas.
Figure 1: Example ML lifecycle with Atlas. Striped boxes and bolded arrows represent Atlas components and data flows. MLaaS providers obtain model artifacts (datasets or models) and perform transformations on them. Atlas attestation clients generate and submit digitally signed artifact and transformation metadata to a transparency log. Model users commonly delegate model provenance checking to an Atlas verifier.
4.1 Transparency Service
To provide transparency across stages of the ML lifecycle, MLaaS providers integrate an Atlas attestation client in their system. The transparency log makes all attestation clients’ collected metadata available to verifiers. The four core techniques underlying the Atlas transparency service are designed to be general, allowing them to remain agnostic to the particular ML lifecycle stage or pipeline they are applied to (R6).
4.1.1 Artifact Measurements
The Atlas attestation client within an MLaaS provider cryptographically measures every artifact ingested into and output by an ML pipeline. That is, we use a collision resistant cryptographic hash function to compute unique, immutable measurement values for every artifact. Because any changes to a given artifact result in a different hash value, Atlas verifiers wishing to check whether a given model artifact matches their expectation are able to detect tampering between stages of the lifecyle (R1).
4.1.2 Model Transformation Integrity
In addition to capturing measurements about all inputs and outputs of an ML pipeline, the Atlas attestation client continuously verifies new inputs and monitors an ML system’s execution to collect detailed information about pipeline state and operations that transform the input artifacts. That is, during data processing, the client tracks dataset modifications and preprocessing operations; during model training, Atlas captures state changes in model weights, hyperparameters, and configurations.
To enhance the integrity of the ML artifacts and system at runtime, Atlas runs ML pipeline code within trusted execution environments (TEEs) (e.g., Intel TDX [48] or AMD SEV-SNP [49]), which serve as a hardware-based root of trust for ML system component measurements extending from its hardware to the firmware and software.
These measures allow the Atlas attestation client to verify the integrity and authenticity of the compute environment before beginning the execution of the pipeline. Throughout execution, the TEE maintains isolated memory regions, reducing the risk of interference with the pipeline, as well as unauthorized disclosure of artifacts and pipeline code, by a compromised MLaaS provider (R4 & R5).
At the conclusion of an operation in a given pipeline, the client generates a record with all collected ML system information, and digitally signs it with a TEE-generated key along with all artifact measurements, providing a model transformation attestation (R2). If required by the artifact producer, Atlas may encrypt the produced model artifacts prior to digitally signing and uploading them to a given hub to further enhance artifact authenticity and confidentiality while in transit and at rest (R1 & R5).
4.1.3 Provenance Chains
To allow verifiers to trace an ML model throughout all of its lifecycle stages from data preparation to deployment, Atlas embeds the cryptographic hash of the transformation attestation for every pipeline input that passed verification into the output attestation. These hash values are digitally signed as part of the transformation attestation. Thus, Atlas attestation clients establish an authenticated, cryptographically verifiable provenance chain representing a model’s lineage relationships throughout its lifecycle (R3).
4.1.4 Transparency Log
On the server side, a transparency log contains the golden values (i.e., known good values) of a producer’s model artifacts, an MLaaS provider’s system components, and the transformation attestations collected by the clients throughout the ML lifecycle. To enable efficient insertion and provenance verification while accommodating the cyclical nature of the ML lifecycle, Atlas relies on two data structures.
First, to provide cryptographic tamper-evidence for the stored values, the transparency log is constructed using an append-only Merkle tree [50], meaning that pipeline metadata can be efficiently inserted in the right-most empty leaf node of the tree (e.g., as in [28]). Second, to enable more efficient verification of provenance across pipelines (or even cycles of the ML lifecycle), Atlas can represent each discreet pipeline/cycle using a different Merkle tree and chain the trees by embedding the Merkle root hash of the upstream pipeline or previous cycle into the “current” Merkle tree’s root (e.g., as in [51]), limiting tree traversal operations when verifying a lifecycle over time (R7).
We note that the choice of whether and how to partition the transparency log’s Merkle trees will result in different security and performance tradeoffs; depending on the party that operates the transparency service in practice, they may favor maintaining a smaller or larger number of trees according to their specific metadata access control and resource requirements.
4.2 Verification System
The Atlas verification system performs staged verification. When a pipeline ingests an input artifact, the attestation client requests the artifact’s verification at the verification service, interacting with the transparency log to obtain the golden values and provenance information.
First, the verification service validates that the received metadata was produced by the expected parties via their digital signatures, and checks whether the artifact matches the expected golden value measurements. If the first verification stage passes, Atlas inspects the received transformation metadata to confirm the recorded ML system and pipeline operations ran as expected based on their TEE-backed attestations. Third, the artifact’s lineage is validated by tracing its provenance chain to check its transformations as far back as specified in the verifier’s policy.
To enable efficient batch verification of input artifacts, Atlas groups related artifacts including training datasets, model weights, algorithm implementations, and evaluation results. By maintaining a cache of verified states and an incremental verification flow, the system avoids re-computing cryptographic operations for unchanged artifacts (R7). This particularly benefits ML pipelines where only small portions of the workflow change between iterations.
5 Implementation
Our proof-of-concept implementation integrates with Kubeflow [33] through standard APIs for metadata tracking and execution monitoring within ML pipelines. This integration approach enabled us to avoid significant modifications to our case study pipelines (§6), while maintaining Atlas’ security and transparency enhancements.
The Atlas attestation client is implemented as two components. First, the metadata sidecar (§5) captures ML artifact and pipeline metadata in C2PA manifest format [10]. Due to current poor support for automated C2PA manifest generation and verification for ML models, we implemented Rust-based library in 2.3k LoC that captures cryptographic measurements, TDX attestation hashes, digital signatures, and temporal metadata. Second, the attestation client implementation leverages a dedicated TEE service (§5, which interacts with Intel TDX [48], a virtualization-layer TEE, to provide the hardware-based security primitives for Atlas.
To implement the transparency log, our framework extends the Sigstore’s Rekor  [28] tamper-evident ledger. Building on Rekor’s existing capabilities, we added support for storage and verification of Atlas C2PA-based model transformation attestations. The modified implementation validates C2PA signatures and attestation measurements during manifest submission, ensuring only properly signed and attested artifacts are stored in the log.
For ease of implementation, our prototype verification system (§5.3) is integrated into the client’s metadata sidecar.
Metadata Sidecar. To integrate our C2PA tooling non-intrusively with ML pipelines, the Atlas attestation client uses a sidecar container architecture. This allows the client to monitor the pipeline for modifications, generate manifests and update relationships through an optimized storage layer. In addition to collecting C2PA manifest metadata, the sidecar generates the authenticated transformation attestations for model artifacts, and tracks artifact relationships by mapping datasets to checkpoints, checkpoints to models, and maintaining configuration bindings.
The sidecar container implementation monitors the pipeline for modifications, generating manifests that are first stored in a local cache layer before being committed to the transparency log. The local cache maintains an indexed hierarchy of manifests for efficient validation during pipeline execution, before the final storage in Rekor for tamper-evident provenance tracking. For storage optimization, our implementation decomposes manifests into constituent components. The assertion store, claim signatures, and metadata are stored separately, with relationships maintained through a reference system. This approach enables efficient updates to specific manifest components, reduced storage redundancy, optimized query performance, and scalable version tracking.
TEE Service. To obtain the Intel TDX-based compute environment attestations that are included in the C2PA metadata, the sidecar interfaces with the TEE service. We implement a dedicated RESTful service that captures TEE state and validates runtime ML system component measurements, which are cryptographically anchored in hardware via Intel TDX. The TEE service implements the Confidential Containers (CoCo) Attestation format [52] for standardized claims used in TEE verification.
We optimize attestation operations through batching and caching mechanisms.The client coordinates with the verification system through standard Kubeflow interfaces, enabling seamless integration with the other Atlas components while preserving attestation integrity.
5.1 Kubeflow Integration
The integration with Kubeflow is achieved through custom operators and controllers that monitor pipeline execution through Kubeflow’s Metadata V2 Beta API and KFP API. Through the //apis/v2beta1/metadata endpoint, we track execution contexts and maintain verifiable records of pipeline runs.
By interfacing with /apis/v2beta1/artifacts, we track model artifacts and their lineage. The metadata store provides structured information about component dependencies and data flow through the /apis/v2beta1/connections endpoint. Our system correlates this information with integrity measurements and hardware attestations, creating verifiable records of pipeline execution states.
The metadata extraction leverages Kubeflow’s event system through /apis/v2beta1/events, enabling real-time capture of pipeline state transitions, component execution details, artifact generation events, and parameter updates. This structured approach enables verification of pipeline states while maintaining compatibility with existing workflows.
5.2 TEE Service Claims
For each attestation request, the client generates a JSON Web Token  [53], which adds custom claims to represent Atlas-specific information including issuer identification and public key for signature verification, temporal validity through expiration and effective timestamps, hardware TEE information and software measurements, and golden values used in verification policy enforcement
Verification Service Implementation. The metadata sidecar also serves as a verification endpoint allowing pipeline components to validate artifact integrity against stored attestations. We optimize the performance of our staged verification system in three ways: 1) by processing changes incrementally and caching to avoid re-verifying unchanged components, 2) via batch processing of verification operations, and 3) parallel verification paths for independent component classes.
For provenance chain validation, our implementation parallelizes verification of cross-component dependencies, version compatibility and evaluation result consistency. The system also maintains verification checkpoints that serve as trusted reference points, enabling partial verification from the last known good state instead of complete chain recomputation.
Error handling focuses on computational efficiency through targeted cache invalidation rather than complete cache clearing. When verification failures occur, the system preserves verified states while ensuring security through selective invalidation. Preliminary testing shows these optimizations reduce verification time by up to 50% through parallelization and caching, while maintaining security guarantees.
5.3 Verification Flow
Figure 2: Verification workflow implementation showing how components are classified and processed: Changes trigger incremental validation that preserves cached states for unchanged components. The system groups related artifacts (training data, model weights, configurations) for batch processing, validates their relationships in parallel, and selectively invalidates only affected components on failures. This approach reduces verification overhead while maintaining security guarantees.
5.4 Framework Adaptability
The BERT deployment validated our framework’s flexibility across different ML environments. The abstraction layer successfully handled variations in framework-specific interfaces, from PyTorch’s hook mechanisms to TensorFlow’s Keras callbacks, while maintaining consistent security guarantees. Custom adapters enabled integration without modifying existing ML infrastructure, demonstrating the framework’s ability to enhance security while preserving established workflows.
6 Evaluation
We validate our framework through a security analysis, preliminary performance testing and two case studies: BERT Meta [3, 4] and BGE reranker [5].
6.1 Security Analysis
Atlas provides measures against the threats outlined in §3.3 through multiple security mechanisms.
For MLaaS provider threats, the hardware-rooted TEEs in Atlas isolate sensitive computations and detect malicious insider tampering with executing ML pipelines. The attestation client continuously validates the runtime environment, generating ML system measurements that are cryptographically bound to model artifacts.
Against hub threats, the Atlas verification system ensures artifact integrity through cryptographic measurements and signatures. The provenance chain maintains verifiable links between artifacts, preventing mismatched signatures or compromised dependencies from propagating.
To address artifact producer threats, Atlas relies on comprehensive provenance tracking that captures all dependencies and modifications. The transparency layer provides an immutable record of pipeline operations, enabling detection of undeclared dependencies or intentional omissions.
6.2 Preliminary Performance Analysis
Our implementation addresses the computational overhead of security mechanisms through several optimization strategies. Preliminary tests demonstrate that Atlas’ security mechanisms introduce minimal overhead, with training overhead remaining under 8% across all test cases. The verification processes scale linearly with model size, while our caching strategies (see §5) effectively optimize performance for iterative ML workflows; we observe a reduction of up to 50% of verification latency compared to the unoptimized implementation, and near-constant time for cached component verification.
For large-scale operations, parallel processing capabilities maintain performance at scale through: 1) concurrent verification operations when dependencies allow; 2) cache optimization to reduce repeated verification costs; 3) selective cache invalidation for error handling.
We plan to conduct more extensive performance benchmarking as part of future work.
6.3 Case Studies
BERT was selected for its complex architecture and widespread production use. Our implementation covers the complete lifecycle from pre-trained model fine-tuning through deployment, with isolated execution environments for attention computations and weight updates.
To validate Atlas’ adaptability, we also implement security controls for BGE reranker on Intel Xeon processors. Atlas secures both instruction-based and layerwise reranking configurations, with training data structured as JSON records containing query-positive-negative text triplets. The provenance system tracks model adaptations across both base model variations (Gemma-2B and MiniCPM-2B), maintaining verifiable records of hyperparameters and training progressions.
Both implementations demonstrate performance characteristics consistent with our analysis in  §6.2, while preserving training efficiency and security guarantees.
7 Case Study: BERT Pipeline
Our proof-of-concept implementation demonstrates the framework’s capabilities through securing a BERT model development and deployment pipeline. We chose BERT Meta [3, 4] for this case study due to its complex architecture, widespread use in production environments, and frequent fine-tuning requirements that test our framework’s adaptability.
The implementation covers the complete model lifecycle from pre-trained model adaptation through deployment. For BERT’s transformer architecture, we implemented isolated execution environments for attention computations and weight updates. The hardware-backed attestation system provides continuous runtime verification during fine-tuning operations, while our provenance system maintains verifiable records of all model evolution steps from the base model through task-specific adaptations.
Performance analysis shows practical viability with training overhead under 8%, primarily from manifest generation, and verification processes scaling linearly with model size. The system’s parallel processing capabilities and cache optimization strategies maintained performance at scale.
8 Discussion and Conclusion
Our implementation and evaluation reveal several key considerations for practical deployment ML pipeline security. Organizations must balance security requirements with operational efficiency, considering factors like verification frequency, attestation depth, and computational overhead. Atlas’ modular design enables customization of these tradeoffs while maintaining core security guarantees.
Current hardware security architectures present integration challenges, particularly around GPU and accelerator attestation support. While Atlas provides secure boundaries for CPU-GPU data transfers, emerging technologies like confidential GPU computing will enable more comprehensive coverage. Integration with legacy HPC systems also requires careful consideration, as traditional job schedulers often lack modern observability capabilities needed for fine-grained monitoring and attestation.
9 Adoption Considerations
9.1 Storage Optimizations
9.1.1 Storage and Scalability
Our analysis suggests opportunities for optimizing manifest storage through decomposed and hybrid architectures. Rather than storing complete manifests as single documents, separating components like assertions, signatures, and metadata could improve efficiency while maintaining security guarantees. Organizations should consider distributed storage strategies that balance immutability requirements with query performance needs.
9.1.2 Decomposed Storage Model
Instead of storing complete manifests as single documents, the system could decompose manifests into their constituent components. The assertion store, claim signatures, and metadata could be stored separately, with relationships maintained through a reference system. This approach would enable more efficient updates and queries of specific manifest components.
The decomposed model could leverage distributed ledger technology (DLT) [54] for critical manifest components while maintaining bulk data in optimized storage systems. This hybrid approach would:
  • Store cryptographic proofs and signatures on the distributed ledger for immutability
  • Maintain manifest metadata and relationships in graph databases for efficient querying
  • Use object storage for large artifacts like model weights and datasets
  • Link components through cryptographic references preserved in the ledger
9.1.3 Hybrid Storage Architecture
A hybrid approach could maintain critical verification data in Rekor for its transparency guarantees while storing detailed manifest data in optimized storage systems. This would balance the need for immutable proof of existence with efficient data access and management.
These storage optimizations could significantly reduce operational overhead while maintaining the security guarantees of our framework. Performance testing indicates potential reduction in storage requirements and query latency through these alternative approaches.
9.2 Deployment Guidelines
9.2.1 Organizational Adaptations
Organization-specific adaptations are necessary to align with existing infrastructure and security policies. Key considerations include:
  • Integration with current MLOps platforms
  • Alignment with existing security monitoring systems
  • Customization of verification policies
  • Adaptation to specific hardware security capabilities
  • Compliance with organizational security standards
9.2.2 Security Requirement Balance
Security requirement balance directly impacts operational efficiency. Organizations must determine appropriate verification frequencies and depth based on their risk profile and performance requirements. For instance, continuous hardware attestation of all pipeline components provides maximum security but introduces significant overhead. A more balanced approach might implement full verification at critical pipeline stages while using lightweight checks during intermediate steps.
9.2.3 Computational Overhead Management
Computational overhead management becomes crucial when scaling the framework across large ML operations. Our implementation shows that intelligent caching of verification results and batch processing of integrity checks can significantly reduce overhead. Organizations should consider:
  • Strategic placement of verification checkpoints - Organizations can tailor verification intensity based on their specific security needs and operational context. While financial or healthcare institutions might require comprehensive verification throughout their ML pipeline, research or development environments might focus verification efforts primarily on model publication or deployment stages. This flexible approach enables efficient resource utilization while maintaining appropriate security levels for each use case.
  • Optimization of hardware attestation frequency - By analyzing pipeline characteristics and risk patterns, attestation frequency can be tuned to concentrate on high-risk operations while reducing overhead during stable processing phases.
  • Efficient manifest storage and retrieval mechanisms - The system maintains an indexed store of manifests with hierarchical organization, enabling quick validation of model lineage while managing storage overhead for long-term provenance tracking.
  • Parallel verification processing where possible - This approach utilizes available computational resources effectively by running verification operations concurrently when component dependencies allow.
10 Future Work
Several potential enhancements could extend our framework’s capabilities and applicability:
Distributed Training Support. The current framework could be enhanced to handle multiple TEEs coordinating across training nodes, with cross-node attestation and verification protocols. This would require developing protocols for maintaining integrity across distributed components while managing the additional complexity of verifying inter-node communications and state synchronization [55].
Federated Learning Compatibility. Our current framework could be extended to support federated learning environments, particularly through integration with Intel’s OpenFL (Open Federated Learning) [56] framework. OpenFL’s architecture, which separates aggregator and participant nodes while maintaining model security, presents unique opportunities and challenges for provenance tracking. The framework would need to extend its attestation and verification protocols to handle distributed model updates while preserving the privacy guarantees inherent in federated learning.
Key considerations include tracking model aggregation operations, verifying participant contributions, and maintaining cryptographic proofs across federation rounds. OpenFL’s existing security features, including its support for secure aggregation and TEE integration, provide natural integration points for our provenance framework. The challenge lies in extending our verification protocols to handle the partial model updates and differential privacy mechanisms [57] common in federated learning scenarios.
Enhanced Scalability Features. Support for more complex ML architectures, particularly for multi-model systems and ensemble methods, would expand the framework’s utility. This would require developing verification protocols for model composition and interaction, tracking dependencies between component models, and maintaining provenance across model combinations[58].
The framework could also be extended to support dynamic trust models, allowing for flexible trust relationships between different components and participants in the ML pipeline.
Algorithmic Security Enhancements. Our framework’s modular design allows for integration of additional algorithmic security methods to enhance pipeline protection. Model watermarking techniques [59, 60] could be incorporated to embed verifiable ownership proofs directly into model weights, providing an additional layer of provenance verification. These watermarks would be included in the manifest chain, creating cryptographically verifiable links between model versions and their origins.
Verification Protocol Extensions. Our staged verification system could be extended to support dynamic trust models. This would allow more flexible verification policies based on component criticality and risk levels, while maintaining our core security guarantees. The current implementation’s classification system provides a foundation for such policy-based verification.
Neural fingerprinting [61, 59] methods could extend our verification capabilities by enabling detection of unauthorized model modifications or derivatives. By maintaining fingerprint signatures in our provenance records, the framework could track model lineage even when traditional hash-based verification is insufficient. This is particularly valuable for scenarios involving fine-tuning or transfer learning.
Property attestation mechanisms could verify specific algorithmic characteristics of models throughout the pipeline. For example, robustness guarantees [62], fairness metrics [63], or backdoor resistance [64] could be measured and included in the manifest chain. These properties would enhance the framework’s ability to detect subtle manipulations that might not affect model hashes but could impact model behavior.
Our work demonstrates that robust security measures can be integrated into ML lifecycles without significant performance impact. The combination of hardware-backed security guarantees with provenance tracking provides a foundation for securing ML pipelines. Our case studies show Atlas’ ability to maintain strong security properties while preserving established development practices.
We identify three directions for further enhancing ML pipeline security: 1. extending hardware security coverage to specialized ML accelerators will enable more comprehensive protection of ML computations; 2. standardizing provenance tracking across the industry will significantly improve interoperability and adoption; 3. developing robust algorithmic verification methods and model guardrails will strengthen protection against sophisticated attacks targeting model behavior, allowing for flexible integration of emerging defenses.

### v2

@@
 Abstract
-The rapid adoption of open source machine learning (ML) datasets and models exposes today’s AI applications to critical risks like data poisoning and supply chain attacks across the ML lifecycle.
-With growing regulatory pressure to address these issues through greater transparency, ML model vendors face challenges balancing these requirements against confidentiality for data and intellectual property needs.
-We propose Atlas, a framework that enables fully attestable ML pipelines.
-Atlas leverages open specifications for data and software supply chain provenance to collect verifiable records of model artifact authenticity and end-to-end lineage metadata.
-Atlas combines trusted hardware and transparency logs to enhance metadata integrity, preserve data confidentiality, and limit unauthorized access during ML pipeline operations, from training through deployment.
-Our prototype implementation of Atlas integrates several open-source tools to build an ML lifecycle transparency system, and assess the practicality of Atlas through two case study ML pipelines.
+We propose Atlas, a framework that enables fully attestable ML pipelines to address ML supply chain risks.
+Atlas leverages runtime pipeline monitoring and open specifications for data and software provenance to collect model artifact integrity and end-to-end lineage metadata.
+Atlas combines trusted hardware and transparency logs to enhance metadata integrity and enable efficient verification of ML pipeline operations, from training through deployment.
+Our prototype implementation of Atlas integrates open-source tools to build an ML lifecycle transparency framework.
 1 Introduction
@@
 The pervasive use of large language models (LLMs), in particular, and multi-stakeholder involvement in model creation and deployment exacerbate security and privacy risks.
-These considerations are emphasized by the global nature and the complexity of large-scale ML deployments with different lifecycle stages:
-  1. 1)
-    Collection and sanitation of a training dataset from several public and proprietary sources.
-  2. 2)
-    Provisioning of the training environment (hardware and software).
-  3. 3)
-    Execution of training across many data centers.
-  4. 4)
-    Construction of a testing dataset from several sources, and the evaluation.
-  5. 5)
-    Deployment and use of the model for inference that is compliant with local laws or corporate policies.
-Each of these stages is vulnerable to malicious or dishonest parties.
+These considerations are emphasized by the global nature and the complexity of large-scale ML deployments with different lifecycle stages (e.g., training dataset collection, execution of training).
+Each stage is vulnerable to malicious or dishonest parties.
 For example, data can be poisoned [1, 2] during collection or training.
 Service providers executing outsourced training can shorten or omit critical steps to reduce their cost.
-Model providers can serve smaller models in SaaS, or even distribute malicious ones.
-On the other hand, responsible model builders and other stakeholders may be incentivised or required to provide security and trust guarantees.
-They may want to prove low bias in their training data, offer easily verifiable performance claims, or guarantee end-to-end integrity of the model creation in high risk domains.
-To address these challenges, it is necessary to guarantee the integrity of the entire ML lifecycle – beginning with the data, through the training, and finally, the evaluation and deployment.
+Popular model hubs hosting pre-trained models are vulnerable to compromises that may result in corrupted, reduced, or malicious model distributions [3, 4]
+On the other hand, recent regulations [5, 6] require model builders and other stakeholders to provide evidence of ML model security and trust.
+They may need to prove low bias in their training data, offer easily verifiable performance claims, or demonstrate end-to-end integrity of model creation in high risk domains.
+To address these challenges, integrity of the entire ML lifecycle must be recorded verifiably – beginning with the data, through the training, and finally, the evaluation and deployment.
 Was the data modified?
@@
 These are example questions that showcase the breadth of the involved challenges that must be tackled to provide end-to-end security.
-In this work, we introduce Atlas, a framework for enhancing the security and transparency of the lifecycle of ML models.
+We introduce Atlas, a framework for enhancing the security and transparency of the lifecycle of ML models.
 Atlas establishes the baseline of fundamental components and capabilities needed for comprehensive provenance tracking at each stage of the ML lifecycle.
-Subsequently, Atlas defines the core integrity requirements for verifiable ML lifecycle transparency.
-We provide a reference implementation that instantiates Atlas using hardware-based security mechanisms – with trusted execution environment (TEE), including attestations.
+Thus, rather than preventing attacks entirely, Atlas detects tampering by verifying the ML lifecycle.
+Atlas addresses two challenges unique to ML lifecycle transparency.
+First, in contrast to the clear dependency trees of traditional software, datasets and algorithm code are tightly coupled in ML models, creating significantly more intricate provenance graphs [3].
+Attesting the integrity of these relationships requires mechanisms that can track cross-organizational transformations across heterogeneous artifacts of varying sizes and formats [7].
+Hence, Atlas monitors ML pipelines during execution and automatically collects ML system information using several data and software provenance frameworks.
+To strengthen the integrity of provenance generation, Atlas relies on hardware trusted execution environments.
+Second, models are not static binaries – they can be further customized by downstream users.
+That is, ML models undergo a non-linear lifecycle where deployment results often necessitate refinement, creating feedback loops between inferencing and data processing [8].
+This iterative process requires special adaptations to provenance tracking mechanisms. Atlas represents these non-linear development paths and enables cryptographic auditing using Merkle trees [9].
 We claim the following contributions:
@@
   2. 2.
-    We instantiate Atlas using TEEs and metadata-based provenance tracking.
+    We instantiate Atlas using Intel Trust Domain eXtensions [10] and metadata-based provenance tracking.
   3. 3.
-    We evaluate our Atlas prototype through two case studies: 1) fine-tuning of a BERT model [3, 4]; 2) fine-tuning of a bge-reranker model [5] .
+    We evaluate our Atlas prototype through a fine-tuning case study with a BERT model [11, 12].
 2 Background & Related Work
-2.1 Data Provenance & Authenticity
-While not a new concern, the authenticity, provenance and attribution of media has received renewed attention in recent years due to the wide online proliferation of images, videos and text that have been manipulated or forged using generative ML models [6, 7].
-One common approach for addressing digital image and video manipulation involves cryptographic techniques such as hashing, digital signatures and blockchains, which enable provenance authentication, tamper detection, and checking for divergent views of the content [8].
-Most notably, the Coalition for Content Provenance and Authenticity (C2PA) specification [9, 10] introduced by the Content Authenticity Initiative (CAI) [11] and Project Origin [12] provides cryptographic binding of authenticity claims, verifiable assertions about content origin and modifications, and tamper-evident audit trails [13].
-Initially launched as a response to the growing challenge of deepfakes  [11] and digital content manipulation, the C2PA standard first gained traction in digital photography and journalism workflows.
-The standard’s extensible and interoperable metadata format has made C2PA an attractive option for ML dataset and model provenance tracking [14].
-Other prior work in this space builds upon digitally signed metadata by leveraging distributed ledger technologies to create transparent and immutable content or provenance records.
-Examples of this approach include AMP [7], the News Provenance Project [15] and Collomosse et al.
-[14].
-Atlas borrows techniques from C2PA and similar approaches, but integrates them directly into the ML system to track model artifact provenance closer to where the transformations occur to achieve higher metadata fidelity, rather than applying them in an ad hoc fashion after a particular dataset or model artifact has been created.
-The Open Source Security Foundation (OpenSSF) Model Signing project [16] is a parallel effort in this space focusing specifically on cryptographic signing of model artifacts to ensure their authenticity and integrity.
-As such, it serves as a crucial building block but requires integration with other tools for more comprehensive supply chain security.
-While Model Signing might integrate with Atlas, additional mechanisms would be needed to address our goal of tracking full model lineage.
+Data Provenance & Authenticity.
+Provenance and attribution of media has recently received attention due to the online proliferation of manipulated or forged content using generative ML models [13, 14].
+Prior work relies primarily on cryptographic hashing and digital signatures to provide data authenticity and integrity.
+The Coalition for Content Provenance and Authenticity (C2PA) specification [15, 16] digitally signs assertions about content origin to provide tamper-evident data audit trails [17, 18].
+C2PA’s extensible metadata format also makes it suitable for ML model artifacts [19].
+The Open Source Security Foundation (OpenSSF) Model Signing project [20] is a parallel effort focusing on integrity and authenticity of trained models.
+Other prior work [14, 21, 22] builds upon hashing and signing with distributed ledger technologies to create transparent and immutable content or provenance records.
+These techniques are crucial building blocks for verifying ML model artifact authenticity and provenance, but each alone is insufficient for end-to-end ML supply chain transparency.
+In contrast, Atlas aims to integrate such techniques into ML systems to track model artifact provenance directly where the transformations occur.
+Supply Chain Integrity.
+Recent cybersecurity regulations [23, 24] have shifted industry focus toward detecting supply chain threats via software dependency tracking with SBOMs [25].
+Similarly, the AIBOM framework [26, 27] focuses on ML model supply chain management.
+Complementing BOM, efforts like OpenSSF Supply Chain Levels for Software Artifacts (SLSA) [28] and SPDX Build [29] collect build provenance, i.e., metadata describing how a particular artifact was produced.
+This approach is also being considered for ML model fine tuning [30].
+Building on such supply chain metadata efforts, a number of frameworks [31, 32, 33] provide mechanisms for collecting, digitally signing and verifying authenticated claims across supply chain steps.
+Atlas borrows concepts from supply chain integrity to support multiple types of software artifact provenance at any stage of the ML lifecyle, providing a more comprehensive view of a model’s supply chain.
+Works about evidence for other properties of the ML lifecycle such as assurance [8] are complementary.
+Model Lineage Tracking.
+The EQTY Lineage Explorer [34] tracks model artifacts throughout the training process, capturing relationships between datasets, model checkpoints and hyperparameters.
+However, unlike Atlas, it lacks cryptographic authenticity properties and focuses primarily on manually collected development-time lineage, rather than automatically capturing and linking information across the entire ML lifecycle.
+ML experiment trackers like Weights and Biases [35], Neptune [36] and Kubeflow Pipelines [37] offer detailed run-time logging of model metadata about training runs, metrics, and model artifacts.
+These tools do not integrate transparently with common ML frameworks, and they typically provide only unauthenticated metadata.
+Atlas, on the other hand, seeks to make model lineage verifiable and support integration into ML frameworks like PyTorch [38].
+Hardware-Based Security for ML.
+Recent developments in trusted execution environment (TEEs) technologies have made it more practical to run large-scale systems and workloads [39, 40], including ML pipelines.
+Chrapek et al.
+[41] deployed and optimized a large language model (LLM) inside a TEE, showing how secure enclaves help protect LLM code and data while in use.
+They maintain practical performance in two TEE configurations based on Intel Software Guard eXtensions (Intel SGX) [42] and Intel Trust Domain eXtensions (Intel TDX) [10].
+Laminator [43] and PraaS [44] demonstrate the application of TEEs to ML model or dataset property attestation and verification.
+Several efforts [7, 45, 46] use TEEs to build confidentiality frameworks for different ML lifecycle stages.
+These works are complementary to Atlas and may enable us to extend our framework.
+Mo et al.’s survey [47] evaluates 38 works that use various TEE implementations to enhance the privacy and integrity of ML training and inference operations.
+The survey highlights several gaps, including the protection of full ML lifecyles, which is the primary focus of Atlas.
+We provide additional background in Appendix A.
+3 System Overview & Threat Model
+3.1 Terminology
+In Atlas, an ML model is composed of several artifacts that include the training dataset, ML algorithm, ML framework (e.g., PyTorch), model configuration (e.g., hyperparameters, weights), and metadata (e.g., license).
+The ML lifecycle consists of various stages, including data preparation, training, evaluation and deployment.
+A common synonym for ML lifecycle is “ML supply chain”, so we use these terms interchangeably in the paper.
+The ML pipeline defines the sequence of operations or a workflow that transforms a model artifact at a particular stage of the ML lifecycle [48].
+To support standardization and repeatability, the pipeline also facilitates workflow management and automation.
+The ML system is the set of hardware and software components that implement and execute an ML pipeline.
+For example, an ML system for training may include orchestration tools, an authentication service, storage systems, automation infrastructure, and specialized compute hardware (e.g., GPUs, TPUs, or custom accelerators).
+In Atlas, metadata describes two mains aspects about an ML model.
+First, provenance metadata refers to the origin and history of custody of a model artifact, including its history of transformations as it traverses the ML lifecycle.
+Pipeline metadata, in turn, describes the ML systems and specifics about the operations that produced a model artifact.
+An attestation in Atlas refers to any digitally signed metadata and serves as evidence for model artifact or ML system authenticity, integrity and provenance.
+Attestations may be generated by hardware or software.
+3.2 System Model
+In Atlas, we target the multi-stakeholder ML model lifecycle setting.
+3.2.1 Stakeholders
+The Artifact Producers are individuals and organizations that create ML model artifacts to provide or sell them to other parties.
+Since the artifacts may represent intellectual property and/or make use of personally identifying information (PII), producers have business and regulatory reasons to preserve their confidentiality.
+To save costs, artifact producers often outsource the operation of ML systems to external service providers.
+ML-as-a-Service (MLaaS) providers operate and maintain the compute infrastructure needed to run ML systems.
+MLaaS providers may offer ML-specific services that leverage general-purpose compute (e.g., [49, 50, 51]), or provide special-purpose systems (e.g., [37, 52]) that can build and run third-party ML systems.
+A Hub is a system that stores and distributes model artifacts.
+Thus, model pipelines typically ingest and output artifacts to and from hubs during their execution, enabled by interfaces exposed by MLaaS providers.
+Hubs may be operated by artifact producers themselves or by third-party vendors, containing open or closed source artifacts (e.g., [53, 54]).
+A Transparency Service in Atlas is responsible for generating, storing and distributing the metadata necessary to verify the authenticity, integrity and provenance of model artifacts.
+We envision model vendors and independent parties operating transparency services in practice.
+Transparency services interface with MLaaS providers through attestation clients that run alongside ML systems to obtain and attest provenance and pipeline metadata.
+On the server side, a transparency log contains the known good values (i.e., golden values) of model artifacts and ML system components, submitted by model producers and MLaaS providers, as well as attestations collected by the clients throughout the ML lifecycle.
+Verification Services evaluate or audit a particular ML model’s lifecycle with the goal of detecting unintended or malicious tampering with the model at any stage.
+In practice, model users, vendors or regulatory entities may operate verification services.
+Using the golden values111 Golden values should be independently verified to establish their trustworthiness.
+Current approaches for auditing golden values include reproducibility [55] and endorsements [56].
+Atlas is agnostic to the chosen method and assumes that evidence of golden value verification can be made available via a transparency service.
+and attestations obtained from a transparency service, a verification service evaluates each ML pipeline and artifact of interest against a set of expectations.
+For example, a model producer may check that the MLaaS provider ran the expected pipeline code, or a model user may verify that a fine-tuned model was produced from the expected foundation model.
+A Model User interacts with a model in an inferencing pipeline, or in a downstream ML pipeline as a dependency, such as a fine-tuning or evaluation pipeline (see §3.2.2).
+3.2.2 ML Lifecycle
+In Atlas, we consider four high-level stages in the ML lifecycle.
+Each builds upon the outputs and feedback from the others, forming a continuous cycle in which models evolve based on real-world usage.
+1. Data processing: Raw data is collected, sanitized and processed into smaller units (e.g., tokens) and collated into a structure ingestable during training or evaluation.
+2. Training: A training algorithm processes a given dataset using an ML system. The output is an ML model.
+3. Evaluation: Following training, model properties like its performance and accuracy undergo further fine-tuning and evaluation using a testing dataset.
+4.
+Deployment: After training and evaluation, an ML model is deployed to a production system configured for inferencing.
+New data obtained from clients during inference are sent back to a data processing pipeline to enhance the training dataset and the model.
+Model use must comply with local laws or corporate policies.
+3.3 Threat Model
+We consider an adversary whose goal is to produce a tampered artifact, e.g., containing a hidden malicious component, so that a transparency service generates a legitimate signature on the artifact or its metadata.
+Thus, Atlas aims to detect such tampering introduced via the ML supply chain.222 Analogously to software correctness (which also applies to ML algorithms and systems), establishing dataset benignity, model quality and safety is a complementary area of research that relies on certifications, e.g., adversarial robustness [57], differential privacy [58], or poisoning [59].
+We leave extending Atlas with such mechanisms as future work.
+Specifically, we consider tampering by MLaaS providers, hubs and artifact producers, while model users, transparency and verification services are trusted in Atlas.
+Compromised MLaaS providers and hubs may involve malicious insiders, or external adversaries seeking to subvert these systems by exploiting vulnerable components.
+Given their central position in the lifecycle, MLaaS providers and hubs may thus be able to compromise model integrity at various stages.
+For example, a malicious MLaaS provider can poison the training data during the curation step of the data processing stage leading to backdoors.
+A compromised hub may, for instance, present a dataset or model with a mismatched signature (e.g., to a different model, or any of its component artifacts) to a model user or MLaaS provider, so that pipelines in subsequent stages of the ML lifecycle may ingest compromised dependencies.
+As a result, these compromises propagate through the ML lifecycle if they go undetected, ultimately leading to vulnerable ML models at the deployment stage.
+This risk is exacerbated if a hub colludes with an MLaaS provider to introduce or accept compromised ML pipeline inputs.
+Artifact producers, on the other hand, may seek to compromise ML models to bypass regulatory requirements, introduce exploitable vulnerabilities or steal private information for profit (e.g, [60]).
+Thus, producers may collude with other untrusted stakeholders, or intentionally inaccurately declare their dependencies, to undermine the integrity of their artifacts.
+3.3.1 Trusted Parties
+Atlas considers the model users, transparency and verification services in an ML lifecycle to be trusted.
+We make the following assumptions about the systems supporting these stakeholders: 1) the hardware and cryptographic primitives are implemented correctly and do not contain known vulnerabilities; 2) a separate PKI system exists and organizations representing the stakeholders follow best practices for key management, network security and access control.
+Further, running attestation clients in TEEs allows us to trust Atlas metadata generation, or detect attempts of tampering by malicious MLaaS providers.
+Similarly, model users and verification services can trust the integrity of golden values and attestations stored in Atlas transparency logs (or detect tampering) via their tamper-evident construction (see §4.2).
+Verification services are trusted to properly evaluate attestations, including their digital signatures, against pre-specified model user expectations.
+3.3.2 Out of Scope
+Many available TEEs provide confidentiality features, but addressing PII and model intellectual property concerns end-to-end requires a more comprehensive framework (e.g.,  [61, 7]).
+We plan to explore confidentiality within Atlas as future work.
+While a critical threat to deployed AI applications, Atlas does not address inference time black-box attacks (e.g., evasion attacks, model extraction, membership inference) caused by malicious users; solutions to reduce this risk [62, 63, 64] are complementary.
+Side-channel attacks against hardware enclaves, physical attacks on hardware infrastructure, and network-level denial of service attacks are also beyond the scope of Atlas.
+These attacks are the subject of a large body of prior work [47], and these complementary security measures could be added to deployments of Atlas.
+3.4 Design Requirements
+We define the following integrity and operational requirements for Atlas:
+R1: Artifact tampering is detectable. To provide model artifact integrity, Atlas must enable verification services to detect unexpected modifications to model artifacts.
+R2: Every model transformation is attested.
+Because adversaries may seek to tamper with model artifacts after they are produced by a pipeline, Atlas attestation clients must record every model transformation in authenticated metadata as evidence for the process, including all inputs to the transformation.
+R3: Verifiable model lineage.
+Atlas verification services must be able to detect unintended/malicious changes by MLaaS providers to the expected stages of the lifecycle (e.g., pipelines operating out of order, or being omitted), from initial data processing through model deployment.
+R4: Strongly isolated ML systems.
+To detect tampering with a pipeline during its execution, Atlas must restrict access to its ML system by malicious MLaaS providers, and contain compromises from propagating beyond the execution environment.
+R5: Pipeline agnostic. To facilitate adoption, Atlas must be agnostic to any ML pipeline that integrates it.
+R6: Efficiency. We seek to minimize the computational and storage overheads incurred by Atlas to enable the implementation and deployment of Atlas in ML systems using commodity platforms and services.
+4 Atlas Framework
+Atlas introduces two core components to the ML lifecycle: 1) the transparency service interacting with MLaaS providers; 2) the verification service for validating model integrity and provenance.
+The core techniques underlying the transparency and verification services are designed to be general, allowing them to remain agnostic to the particular ML lifecycle stage or pipeline they are applied to (R5).
+Fig.
+1 depicts an example ML model lifecycle with Atlas.
+Figure 1: Atlas workflow for ML lifecycle transparency in a BERT Meta [11, 12] fine-tuning process.
+The attestation client monitors ML systems running in TEEs throughout the pipeline stages, collecting provenance metadata from initial deployment through verification.
+4.1 Attestation Client
+Atlas combines mechanisms for artifact and runtime environment integrity to provide transparency across the stages of the ML lifecycle.
+Thus, MLaaS providers integrate an Atlas attestation client with an ML system, each running within a dedicated trusted execution environment (TEE) (e.g., Intel TDX [10] or AMD SEV-SNP [65]).
+TEEs serve two purposes in Atlas.
+First, TEE hardware-enforced memory integrity detects runtime tampering with the attestation client and ML system components by the MLaaS host (including privileged software like the OS or hypervisor).
+Second, TEEs provide a hardware-based root of trust for provenance and pipeline metadata.
+4.1.1 Artifact Measurements
+For every artifact that is ingested into and output by an ML pipeline, the attestation client computes a cryptographic measurement using a collision resistant hash function resulting in a unique, immutable identifier.
+If an artifact is tampered with, its measurement will differ from its golden value, allowing Atlas verification services to detect modifications between lifecycle stages (R1).
+Artifact producers are expected to publish digitally signed measurements as golden values whenever an artifact is first created.
+4.1.2 Model Transformation Integrity
+The attestation client is also responsible for generating provenance and pipeline metadata describing the transformation process and ML system that produced a new model artifact.
+Before pipeline execution begins, the client obtains a hardware attestation to its initial state from its TEE, which includes measurements of the client’s execution environment.
+At the start of pipeline execution, the attestation client remotely attests the ML system’s TEE [10, 65], verifying the integrity of the pipeline’s compute environment.
+Specifically, the client checks that the ML system’s firmware, OS and pipeline code match the golden values published by the MLaaS provider to ensure that a pipeline starts from a known good state.
+[^3] Because only select TEE implementations [66] provide built-in support for attested interactions with I/O devices or ML accelerators like GPUs, it is challenging to distinguish between benign and malicious runtime modifications to pipelines via the network, disk, etc (see §6.4).
+Throughout pipeline execution, TEE hardware enforces memory integrity checks and isolation of executing attestation client/ML system code and in-memory data, reducing the risks of interference by compromised MLaaS providers or any co-located ML systems (R4).
+Further, the attestation client continuously monitors an ML system’s execution, which allows it to determine when artifacts move into or outside of the system, and to collect information about operations that transform the input artifacts (see §5).
+For example, during data processing, the client tracks dataset modifications and preprocessing operations; during model training, Atlas captures state changes in model weights, hyperparameters, and configurations.
+When pipeline execution concludes, the attestation client generates pipeline metadata containing all collected ML system runtime information and the ML system’s TEE attestation.
+Then, the client creates provenance metadata including 1) artifact measurements, 2) operations producing the outputs, 3) TEE attestation for the client, 4) pipeline metadata.
+The attestation client then digitally signs the provenance metadata with keys it generates within the TEE, cryptographically binding ML artifacts to the pipeline and precursor artifacts that created them in a transformation attestation (R2).
+The client uploads this attestation to the transparency service.
+4.1.3 Provenance Chains
+To enable ML model provenance tracking throughout all of its lifecycle stages, the attestation client embeds the cryptographic hash of precursor artifact attestations into every artifact’s transformation attestation.
+These hash values are digitally signed as part of the transformation attestation, enabling detection of unexpected/malicious modifications between attested ML artifact transformations.
+Thus, Atlas attestation clients establish an authenticated, verifiable provenance chain representing a model’s lineage relationships (R3).
+4.2 Transparency Log
+The transparency service’s log makes all published golden values and attestation client-generated metadata available to verification services.
+To enable efficient insertion and provenance verification while accommodating the cyclical nature of the ML lifecycle, Atlas relies on two data structures (R6).
+First, to provide cryptographic tamper-evidence for the stored values, the transparency log is constructed using an append-only Merkle tree [9], meaning that pipeline metadata can be efficiently inserted in the right-most empty leaf node of the tree (e.g., as in [32]).
+Second, to enable more efficient verification of provenance across pipelines (or even cycles of the ML lifecycle), Atlas can represent each discrete pipeline/cycle using a different Merkle tree.
+These separate trees are linked by embedding the Merkle root hash of the preceding pipeline or cycle into the “latest” Merkle tree structure (e.g., as in [67, 68]), providing a temporal cryptographic tree chain.
+Optimizations to chained Merkle trees have been developed in prior research [68, 69, 70].
+4.3 Verification Service
+Stakeholders in Atlas seek to validate that artifacts have not been tampered with and that they were produced by expected pipelines running in high-integrity execution environments.
+To respond to verification requests, the Atlas verification service obtains golden values and transformation attestations from transparency logs relevant to an artifact in question.
+First, the verification service validates the digital signatures on attestations and golden values to authenticate their producers.
+It then checks whether the artifact matches its golden value.
+If these checks pass, the service inspects the transformation attestations to confirm the ML system and pipeline operations ran as expected based on TEE attestations and golden values.
+Atlas validates artifact lineage by traversing the provenance chain, enabling efficient verification through batching related artifact types and maintaining a cache of verified transformations.
+This avoids repeated inspection of unchanged artifacts (R6), which particularly benefits iterative ML pipelines.
+5 Implementation
+Our proof-of-concept implementation integrates with PyTorch [38] and Kubeflow [37] through standard APIs for metadata tracking and execution monitoring within ML pipelines.
+This integration approach enabled us to avoid significant modifications to our case study pipelines (§6), while maintaining Atlas’ security and transparency enhancements.
+We leverage Intel TDX [10], a virtualization-layer TEE, to provide the hardware-based security primitives for ML systems and attestation clients in Atlas.
+The attestation client is implemented as two components.
+First, a continuous ML system monitor integrates with PyTorch to collect metadata for a given pipeline.
+Second, the metadata sidecar (§5) running inside a dedicated Intel TDX TEE generates ML artifact and pipeline metadata in C2PA manifest format [16].
+Due to current poor support for automated C2PA manifest generation for ML models, we implemented a Rust-based library and CLI [^4] Available at github.com/IntelLabs/atlas-cli that captures artifact measurements, Intel TDX attestations, and digital signatures in C2PA format.
+Supporting other software provenance formats [28, 29] is future work.
+We extend Sigstore’s Rekor [32] to support Atlas C2PA-based model transformation attestations, validating signatures and measurements to ensure only properly signed artifacts are stored.
+Out of space considerations, we provide additional details about the implementation in App. B.
+Atlas Workflow Example. We illustrate Atlas’s end-to-end operations through an example with fine-tuning a BERT model for sentiment analysis:
+  1. 1.
+    Pipeline Environment Provisioning: MLaaS provider sets up Atlas attestation client and ML system monitor in Kubeflow.
+  2. 2.
+    Data Preparation: Data scientist prepares and uploads custom dataset, with Atlas metadata sidecar measuring and attesting the dataset using Intel TDX, submitting attestation to the transparency log.
+  3. 3.
+    Environment Validation: Atlas sidecar verifies training environment integrity, adding TDX-based attestation to the C2PA manifest.
+  4. 4.
+    Training Process Monitoring: Attestation client tracks (App. B.3):
+    • Model weight changes via PyTorch hooks
+    • Checkpoint creation and modifications
+    • Hyperparameter updates
+  5. 5.
+    Model Verification: Third-party regulator verifies model provenance using Atlas verification service.
+  6. 6.
+    Deployment: Model vendor deploys verified model with provenance chain for user and application integrity validation.
+Metadata Sidecar.
+Because the Atlas ML system monitor runs alongside untrusted MLaaS provider code, the attestation client’s metadata sidecar leverages TEE remote attestation to detect tampering with the ML system monitor.
+That is, the sidecar interfaces with the ML system to obtain the Intel TDX-based compute environment attestations that capture TEE state and ML system component measurements, which are cryptographically anchored in hardware.
+We use the Confidential Containers (CoCo) framework [71] to implement the remote attestation procedure in the sidecar and ML system monitor.
+Once the ML system monitor’s integrity has been validated, the sidecar generates and digitally signs C2PA manifests.
+These include the sidecar’s and ML system’s Intel TDX attestations, the received ML system metadata, the measurements for the pipeline’s input and output artifacts, and hashes for any linked transformation attestations.
+We describe a storage optimization in App.
+B.5.
+Verification Service Implementation.
+For ease of implementation, the metadata sidecar also serves as a verification endpoint allowing pipeline components to validate artifact integrity against stored attestations.
+We optimize the performance of our staged verification system in three ways: 1) by processing changes incrementally and caching to avoid re-verifying unchanged components, 2) via batch processing of verification operations, and 3) parallel verification paths for independent component classes.
+App.
+B.6 provides additional details.
+6 Evaluation
+We validate our framework through a security analysis, preliminary performance testing and a case study with a BERT Meta [11, 12] fine-tuning pipeline.
+6.1 Security Analysis
+Atlas provides measures against the threats outlined in §3.3 through multiple security mechanisms.
+For MLaaS provider threats, the hardware-rooted TEEs in Atlas isolate sensitive computations and detect malicious insider tampering with executing ML pipelines.
+The attestation client continuously validates the runtime environment, generating ML system measurements that are cryptographically bound to model artifacts.
+Atlas counters hub threats by verifying artifact integrity through cryptographic measurements and signatures, maintaining a provenance chain that identifies mismatched signatures or tampered dependencies before they propagate.
+Atlas mitigates artifact producer threats through comprehensive provenance tracking, providing an immutable record of pipeline operations that detects undeclared dependencies and intentional omissions.
+6.2 Preliminary Performance Analysis
+We conduct our experiments on Intel® Xeon® Gold 5520+ processors with 256 GB of RAM running Ubuntu 24.04 beta.
+Employing Atlas with the BERT Meta case study’s CPU-only PyTorch-based fine-tuning pipeline, where the provenance chain covers $20$ artifacts (up to $120$ for more complex pipelines).
+Our measurements demonstrate near-linear scalability of verification time across different chain lengths and model sizes.
+Preliminary tests show Atlas security mechanisms introduce minimal training overhead (under 8%), with each C2PA provenance manifest (8KB) containing artifact measurements, TEE attestations, and pipeline metadata.
+Verification processes scale linearly with model size, and our caching strategies reduce verification latency by up to 50%, achieving near-constant time for cached (see §5) component verification.
+For large-scale operations, performance is maintained through concurrent verification operations, cache optimization, and selective invalidation for error handling.
+We plan to conduct more extensive benchmarking in future work.
+6.3 Case Study
+BERT was selected for its complex architecture and widespread production use.
+Our implementation covers the complete lifecycle from pre-trained model fine-tuning through deployment, with Intel TDX TEEs for attention computations and weight updates.
+Atlas secures instruction-based configuration using JSON records with query-positive-negative text triplets, tracking model adaptations and maintaining verifiable records of hyperparameters and training progressions.
+The BERT Meta implementation demonstrates performance consistent with our analysis in §6.2.
+6.4 Discussion & Limitations
+Our implementation reveals limitations in current hardware security.
+While Atlas provides TEE-based protection for CPU operations [10], ML workloads rely on GPUs and accelerators without equivalent security features [72], creating security and trust boundaries between protected and unprotected environments [73].
+Emerging solutions like confidential GPU computing show promise but have performance trade-offs [74, 75].
+Additionally, ML lifecycle transparency creates competing requirements between verification and confidentiality of intellectual property and sensitive data [76], an important challenge that Atlas needs to address in a future version.
+Organizations must balance security requirements with operational efficiency, considering factors like verification frequency, attestation depth, and computational overhead.
+7 Conclusion
+The combination of hardware-backed security with runtime provenance tracking in Atlas provides a foundation for securing ML pipelines.
+Our case study shows Atlas’s ability to integrate into existing ML frameworks with reasonable performance.
+Interesting directions for future work include: 1.
+provenance tracking of ML accelerator-based computations, 2.
+end-to-end ML lifecycle confidentiality, and 3.
+algorithmic verification methods and model guardrails against attacks targeting model behavior.
+Appendix A Background & Related Work
+In addition to the works highlighted in §2, we describe further details and approaches about related work addressing ML lifecycle security and integrity.
+A.1 Data Provenance & Authenticity
+C2PA.
+The Coalition for Content Provenance and Authenticity (C2PA) specification [15, 16] was initially introduced by the Content Authenticity Initiative (CAI) [77] and Project Origin [78] as a response to the growing challenge of deepfakes [77] and digital content manipulation, gaining traction in digital photography and journalism workflows.
+LakeFS.
 LakeFS [22] combines Git-like semantics with concepts from object stores such as S3 to provide a version control system for data, including ML datsets.
@@
 Atlas’ metadata centered approach, on the other hand, enables first- and third-party ML dataset consumers to track changes and check their provenance, even when they may not have direct access to the data.
-2.2 Supply Chain Integrity
-While the ML supply chain is an emerging area of research, software supply chain security and integrity has been the subject of much prior work.
-Thus, several of these approaches are being explored for their applicability to enhancing the integrity of the ML model lifecycle.
+A.2 Supply Chain Integrity
+BOM.
 Bills of Materials (BOM) have been employed to document the list of components of a hardware or software product for over three decades [79].
-Due to recent regulations [19, 20], Software BOM (SBOM) have been the focus of many industry and academic efforts seeking to facilitate tracking software dependencies and other metadata to detect supply chain threats [21].
-The AIBOM framework [22, 23] incorporates concepts from traditional software supply chain management while focusing on intended ML model use cases.
+Software BOM (SBOM) have been the focus of many industry and academic efforts seeking to facilitate tracking software dependencies and other metadata [25], to improve their adoption, and to enhance SBOM integrity and privacy (e.g., [80, 81]).
+Similarly, the AIBOM framework [26, 27] focuses on intended ML model supply chain management.
 Like SBOM, AIBOM provide a mechanism for tracking model software dependencies and maintaining model metadata.
 At the time of writing, we are not aware of any frameworks other than Atlas that utilize any sort of BOM data format to track ML model components.
-Complementing BOM, a number of efforts like OpenSSF Supply Chain Levels for Software Artifacts (SLSA) [24] and SPDX Build [25] have sought to collect build provenance, i.e., metadata describing the process through which a particular artifact was produced.
-With its incremental guidance on implementing integrity measures for software builds and its build provenance data format, the SLSA framework [24] is gaining traction among software vendors, and also beginning to see some use for ML model fine tuning [26].
-Atlas’ configurable metadata collection capabilities would enable us to collect build provenance at any stage of the ML model lifecycle, providing a more comprehensive view.
+Authenticated claims.
 A number of frameworks for capturing and verifying a variety of security claims and metadata about the supply chain have been proposed.
 in-toto [31] collects authenticated claims across supply chain steps, including SBOM and SLSA metadata.
-In particular, in-toto enables software development pipeline owners and downstream artifact consumers to specify end-to-end supply chain policies, and validate that only the expected parties carried out specific steps in the supply chain, and artifacts underwent transformations in the expected order.
-Given recent and upcoming enhancements that further generalize the framework, in-toto may be a suitable option for specifying and verifying end-to-end ML model lifecycle integrity policies in Atlas.
+In particular, in-toto enables software development pipeline owners and downstream artifact consumers to specify end-to-end supply chain policies, and validate that only the expected parties carried out specific steps in the pipeline and artifacts underwent transformations in the expected order.
+Given recent and upcoming enhancements that further generalize the framework, in-toto may be a suitable option for specifying and verifying end-to-end ML model pipeline integrity policies in Atlas.
 Sigstore [32] provides a transparency log-based infrastructure for issuing signing credentials and validating digital signatures on supply chain artifacts and metadata.
-Similarly, Supply Chain Integrity, Transparency and Trust (SCITT) [29] is an architecture for implementing distributed ledger to provide global visibility and auditing for supply chain operations and claims.
-The SCITT architecture also includes confidential computing technologies to help ensure that only authorized parties submit claims to the transparency ledger.
-Atlas builds upon several of these supply chain integrity and transparency approaches, but seeks to support multiple types of artifact provenance, supply chain metadata and integrity verification mechanisms all within a customizable, integrated framework designed for the ML model lifecycle.
-2.3 Model Lineage Tracking
-The EQTY Lineage Explorer [30] provides granular tracking of model artifacts throughout the training process.
-The tool captures relationships between datasets, model checkpoints and hyperparameters, enabling developers to trace model evolution.
-Unlike Atlas, it lacks cryptographic guarantees for artifact authenticity and focuses primarily on manually collected development-time lineage, rather than automatically capturing and linking information across the entire ML lifecycle.
-ML experiment trackers like Weights and Biases [31], Neptune [32] and Kubeflow Pipelines [33] offer detailed run-time logging of model metadata about training runs, metrics, and model artifacts.
-While comprehensive for collaborative ML application development, these tools do not necessarily integrate transparently with common ML frameworks, and they typically provide only unauthenticated metadata that makes it less appropriate for establishing trust in a particular ML pipeline.
-Atlas, on the other hand, focuses making model lineage verifiable and on enabling adoption by supporting integration into existing open ML frameworks.
-2.4 Hardware-Based Security for ML
-Recent developments in trusted execution environments (TEEs) technologies have made it more practical to run large-scale systems and workloads.
-Thus, a growing number of proposals leverage TEEs to harden a variety of ML systems through hardware-enforced isolation and integrity verification mechanisms.
-Chrapek et al.
-[34] use TEEs in a large language model (LLM) fine-tuning system running inside secure enclaves that help protect LLM code and data while in use.
-To this end, they evaluate their LLM system using the process-level Intel®Software Guard eXtensions (Intel®SGX) and virtual machine-level Intel®Trust Domain eXtensions (Intel®TDX) TEEs, showing that they are able to maintain practical performance overhead of less than 10% for models like Llama2 in both TEE configurations.
-Laminator [35] demonstrates the application of TEEs to ML property attestation for ML model regulatory compliance verification.
-Specifically, this work leverages TEEs’ hardware-assisted remote attestation and cryptographic binding capabilities to generate verifiable ML model property cards.
-This work is complementary to Atlas and may enable us to extend our framework to support such ML property attestations.
-Mo et al.’s survey [36] evaluates 38 works that use various TEE implementations to enhance the privacy and integrity of ML training and inference operations.
-The survey highlights several gaps that existing TEE-based frameworks do not address, including the protection of a full ML lifecycle, which is the primary focus of Atlas.
-3 System Overview & Threat Model
-3.1 Terminology
-In Atlas, an ML model is composed of a number of artifacts that include the training dataset, ML algorithm, ML framework (e.g., PyTorch), model configuration (e.g., hyperparameters, weights), and metadata (e.g., license).
-The ML lifecycle consists of various stages, including data preparation, training, evaluation and deployment.
-A common synonym for ML lifecycle is “ML supply chain”, so we use these terms interchangeably in the paper.
-The ML pipeline defines the sequence of operations or a workflow that transforms or operates on a model artifact at a particular stage of the ML lifecycle [37].
-To support the standardization and repeatability of a specific workflow, the pipeline is also used to facilitate workflow management and automation, e.g., training or evaluation pipelines.
-The ML system is the set of hardware and software components that implement and execute an ML pipeline.
-For example, an ML system for training may include orchestration tools, an authentication service, storage systems, automation infrastructure, and specialized compute hardware (e.g., GPUs, TPUs, or custom accelerators).
-3.2 System Model
-In Atlas, we target the multi-stakeholder ML model lifecycle setting.
-3.2.1 Stakeholders
-The Artifact Producers are individuals and organizations that create ML model artifacts to provide or sell them to other parties.
-Since the artifacts may represent intellectual property and/or make use of personally identifying information (PII), producers have business and regulatory reasons to preserve the confidentiality of their artifacts.
-To save costs, artifact producers often outsource the operation of ML systems to external MLaaS providers.
-ML-as-a-Service (MLaaS) providers operate and maintain the compute infrastructure needed to run ML systems on behalf of artifact producers.
-MLaaS providers may offer ML-specific services that leverage general-purpose systems (e.g., [38, 39, 40]), or provide special-purpose systems like orchestrators [33] or CI/CD (e.g.
-[41]) that can be used to build and run third-party ML systems.
-A Hub is a system that stores and distributes model artifacts.
-Thus, model pipelines typically ingest and output artifacts to and from hubs during their execution, enabled by interfaces exposed by MLaaS providers.
-Hubs may be operated by artifact producers themselves or by third-party vendors, containing open or closed source artifacts (e.g.,HuggingFace [42], PyTorch Hub [43]).
-A Transparency Service in Atlas is responsible for generating, storing and distributing the information necessary to verify the authenticity and integrity of model artifacts, ML systems and lifecycles as a whole.
-We envision model vendors and independent parties to operate transparency services in practice.
-Verifiers are entities that seek to evaluate or audit a particular ML model’s lifecycle with the goal of checking whether the model it produced was tampered with in an unintended or malicious way.
-To this end, verifiers in Atlas define a set of expectations about a producer’s model artifacts, MLaaS providers, and their ML system components.
-In practice, model users, vendors or regulatory entities may act as verifiers.
-A Model User interacts with a model in an inferencing pipeline, or in a downstream ML pipeline as a dependency, such as a fine-tuning or evaluation pipeline (see §3.2.2).
-3.2.2 ML Lifecycle
-In Atlas, we consider four common high-level stages in the ML lifecycle.
-Each builds upon the outputs and feedback from the others, forming a continuous cycle that allows models to evolve and adapt based on real-world usage.
-1. Data processing: Raw data is collected, curated, processed into smaller units (e.g., tokens) and collated into a structure ingestable during training.
-2.
-Training: a training algorithm processes a given dataset using an ML framework.
-The output of the training pipeline is an ML model that may undergo further fine tuning and evaluation, prior to being deployed for inferencing.
-3. Evaluation: following training, model properties like its performance and accuracy are checked and further adjusted, prior to production release and deployment.
-4.
-Deployment: Once an ML model has been trained and evaluated, it is deployed to a production system configured for inferencing.
-The new data obtained from clients during inference is sent back to a data processing pipeline to enhance the training dataset, and the model.
-3.3 Threat Model
-Atlas is concerned with addressing risks to ML model artifacts introduced via the ML lifecycle, i.e., attack vectors in third-party MLaaS and hubs across different ML pipelines that lead to compromised ML artifacts.
-Specifically, while transparency services, verifiers and model users are trusted in Atlas, we consider threats by MLaaS providers, hubs and artifact producers.
-The adversary’s goal is to produce a tampered artifact or model, e.g., containing a hidden malicious component, so that a trusted transparency service generates a legitimate signature on the artifact or its metadata.
-Compromised MLaaS providers and hubs may involve malicious insiders, or external adversaries seeking to subvert these systems by exploiting vulnerable components.
-Given their central position, MLaaS providers and hubs may thus be able to compromise model integrity at various stages of the ML lifecycle.
-For example, a malicous MLaaS provider can poison the training data during the curation step of the data processing stage leading to backdoors [2, 1].
-A malicious hub may, for instance, present a dataset or model with a mismatched signature (e.g., to a different model, or any of its component artifacts) to a model user or MLaaS provider, so that pipelines in subsequent stages of the ML lifecycle may ingest compromised dependencies.
-As a result, these compromises may propagate through the ML lifecycle if they go undetected, ultimately leading to vulnerable ML models at the deployment stage.
-This risk is exacerbated if a hub colludes with an MLaaS providers to introduce or accept compromised ML pipeline inputs.
-Artifact producers, on the other hand, may seek to compromise ML models in order to bypass regulatory requirements, introduce exploitable vulnerabilities or steal private information for profit (e.g, [44]).
-Thus, producers may collude with other untrusted participants, or inaccurately declare their dependencies leading to intentional omissions or misrepresentations their artifacts, to surreptitiously undermine the integrity of their artifacts.
-3.3.1 Trusted Parties
-As mentioned, Atlas considers the transparency services, verifiers and model users in an ML lifecycle to be trusted.
-We make the following assumptions about the security of the systems underlying and enabling the participants: 1) the hardware and cryptographic primitives are implemented correctly and do not contain known vulnerabilities; 2) a separate PKI system and organizations representing the participants follow best practices for key management, network security and access control.
-This allows us to trust that transparency services correctly implement Atlas attestation mechanisms, detect tampering with them by a malicious MLaaS provider and maintain accurate golden values of model artifacts.
-Verifiers are trusted to properly evaluate Atlas attestations, including their digital signatures, against pre-specified expectations, and to publish a record of their evaluations.
-3.3.2 Out of Scope
-We consider the following threats to be out of scope for Atlas.
-While a critical threat to deployed AI applications, inference time black-box attacks (e.g., evasion attacks, model extraction, membership inference) are caused by malicious users, which we do not address; solutions proposed to reduce this particular risk [45, 46, 47] are complementary to our work.
-Side-channel attacks against hardware enclaves, physical attacks on hardware infrastructure, and network-level denial of service attacks are also beyond the scope of Atlas.
-These attacks are the subject of a large body of prior work [36], and we acknowledge that these complementary security measures could be added to deployments of Atlas.
-3.4 Design Requirements
-We define the following integrity and operational requirements for Atlas:
-R1: Artifact tampering is detectable. To provide model artifact authenticity, Atlas must enable verifiers to detect unexpected modifications to model artifacts by malicious MLaaS providers or hubs.
-R2: Every model transformation is attested.
-Because adversaries may seek to tamper with model artifacts after they are produced by a pipeline, every model transformation in Atlas must produce an authenticated record as evidence for the process.
-R3: Verifiable model lineage.
-Atlas verifiers must be able to detect unauthorized operators and unintended or malicious changes to the expected stages of the lifecycle (e.g., pipelines operating out of order, or being omitted).
-R4: Strongly isolated ML systems.
-To reduce the risks of tampering with a pipeline during its execution, Atlas must restrict access to its ML system by unauthorized entities, and contain compromises from propagating beyond the execution environment.
-R5: Controlled artifact disclosure.
-Atlas must address data and model producer concerns about PII and intellectual property confidentiality by enabling participants to control which parties and systems are able to access datasets and model components.
-R6: Pipeline agnostic. To facilitate adoption, Atlas must be agnostic to any ML pipeline that integrates it.
-R7: Efficiency. We seek to minimize the computational and storage overheads incurred by Atlas to enable the implementation and deployment of Atlas in ML systems using commodity platforms and services.
-4 Atlas Framework
-Atlas introduces two core components to the ML lifecycle: 1) the transparency service interacting with MLaaS providers; 2) the verification service that model users and verifiers use to verify integrity and provenance.
-Fig.
-1 depicts an example ML model lifecycle with Atlas.
-Figure 1: Example ML lifecycle with Atlas.
-Striped boxes and bolded arrows represent Atlas components and data flows.
-MLaaS providers obtain model artifacts (datasets or models) and perform transformations on them.
-Atlas attestation clients generate and submit digitally signed artifact and transformation metadata to a transparency log.
-Model users commonly delegate model provenance checking to an Atlas verifier.
-4.1 Transparency Service
-To provide transparency across stages of the ML lifecycle, MLaaS providers integrate an Atlas attestation client in their system.
-The transparency log makes all attestation clients’ collected metadata available to verifiers.
-The four core techniques underlying the Atlas transparency service are designed to be general, allowing them to remain agnostic to the particular ML lifecycle stage or pipeline they are applied to (R6).
-4.1.1 Artifact Measurements
-The Atlas attestation client within an MLaaS provider cryptographically measures every artifact ingested into and output by an ML pipeline.
-That is, we use a collision resistant cryptographic hash function to compute unique, immutable measurement values for every artifact.
-Because any changes to a given artifact result in a different hash value, Atlas verifiers wishing to check whether a given model artifact matches their expectation are able to detect tampering between stages of the lifecyle (R1).
-4.1.2 Model Transformation Integrity
-In addition to capturing measurements about all inputs and outputs of an ML pipeline, the Atlas attestation client continuously verifies new inputs and monitors an ML system’s execution to collect detailed information about pipeline state and operations that transform the input artifacts.
-That is, during data processing, the client tracks dataset modifications and preprocessing operations; during model training, Atlas captures state changes in model weights, hyperparameters, and configurations.
-To enhance the integrity of the ML artifacts and system at runtime, Atlas runs ML pipeline code within trusted execution environments (TEEs) (e.g., Intel TDX [48] or AMD SEV-SNP [49]), which serve as a hardware-based root of trust for ML system component measurements extending from its hardware to the firmware and software.
-These measures allow the Atlas attestation client to verify the integrity and authenticity of the compute environment before beginning the execution of the pipeline.
-Throughout execution, the TEE maintains isolated memory regions, reducing the risk of interference with the pipeline, as well as unauthorized disclosure of artifacts and pipeline code, by a compromised MLaaS provider (R4 & R5).
-At the conclusion of an operation in a given pipeline, the client generates a record with all collected ML system information, and digitally signs it with a TEE-generated key along with all artifact measurements, providing a model transformation attestation (R2).
-If required by the artifact producer, Atlas may encrypt the produced model artifacts prior to digitally signing and uploading them to a given hub to further enhance artifact authenticity and confidentiality while in transit and at rest (R1 & R5).
-4.1.3 Provenance Chains
-To allow verifiers to trace an ML model throughout all of its lifecycle stages from data preparation to deployment, Atlas embeds the cryptographic hash of the transformation attestation for every pipeline input that passed verification into the output attestation.
-These hash values are digitally signed as part of the transformation attestation.
-Thus, Atlas attestation clients establish an authenticated, cryptographically verifiable provenance chain representing a model’s lineage relationships throughout its lifecycle (R3).
-4.1.4 Transparency Log
-On the server side, a transparency log contains the golden values (i.e., known good values) of a producer’s model artifacts, an MLaaS provider’s system components, and the transformation attestations collected by the clients throughout the ML lifecycle.
-To enable efficient insertion and provenance verification while accommodating the cyclical nature of the ML lifecycle, Atlas relies on two data structures.
-First, to provide cryptographic tamper-evidence for the stored values, the transparency log is constructed using an append-only Merkle tree [50], meaning that pipeline metadata can be efficiently inserted in the right-most empty leaf node of the tree (e.g., as in [28]).
-Second, to enable more efficient verification of provenance across pipelines (or even cycles of the ML lifecycle), Atlas can represent each discreet pipeline/cycle using a different Merkle tree and chain the trees by embedding the Merkle root hash of the upstream pipeline or previous cycle into the “current” Merkle tree’s root (e.g., as in [51]), limiting tree traversal operations when verifying a lifecycle over time (R7).
-We note that the choice of whether and how to partition the transparency log’s Merkle trees will result in different security and performance tradeoffs; depending on the party that operates the transparency service in practice, they may favor maintaining a smaller or larger number of trees according to their specific metadata access control and resource requirements.
-4.2 Verification System
-The Atlas verification system performs staged verification.
-When a pipeline ingests an input artifact, the attestation client requests the artifact’s verification at the verification service, interacting with the transparency log to obtain the golden values and provenance information.
-First, the verification service validates that the received metadata was produced by the expected parties via their digital signatures, and checks whether the artifact matches the expected golden value measurements.
-If the first verification stage passes, Atlas inspects the received transformation metadata to confirm the recorded ML system and pipeline operations ran as expected based on their TEE-backed attestations.
-Third, the artifact’s lineage is validated by tracing its provenance chain to check its transformations as far back as specified in the verifier’s policy.
-To enable efficient batch verification of input artifacts, Atlas groups related artifacts including training datasets, model weights, algorithm implementations, and evaluation results.
-By maintaining a cache of verified states and an incremental verification flow, the system avoids re-computing cryptographic operations for unchanged artifacts (R7).
-This particularly benefits ML pipelines where only small portions of the workflow change between iterations.
-5 Implementation
-Our proof-of-concept implementation integrates with Kubeflow [33] through standard APIs for metadata tracking and execution monitoring within ML pipelines.
-This integration approach enabled us to avoid significant modifications to our case study pipelines (§6), while maintaining Atlas’ security and transparency enhancements.
-The Atlas attestation client is implemented as two components.
-First, the metadata sidecar (§5) captures ML artifact and pipeline metadata in C2PA manifest format [10].
-Due to current poor support for automated C2PA manifest generation and verification for ML models, we implemented Rust-based library in 2.3k LoC that captures cryptographic measurements, TDX attestation hashes, digital signatures, and temporal metadata.
-Second, the attestation client implementation leverages a dedicated TEE service (§5, which interacts with Intel TDX [48], a virtualization-layer TEE, to provide the hardware-based security primitives for Atlas.
-To implement the transparency log, our framework extends the Sigstore’s Rekor  [28] tamper-evident ledger.
-Building on Rekor’s existing capabilities, we added support for storage and verification of Atlas C2PA-based model transformation attestations.
-The modified implementation validates C2PA signatures and attestation measurements during manifest submission, ensuring only properly signed and attested artifacts are stored in the log.
-For ease of implementation, our prototype verification system (§5.3) is integrated into the client’s metadata sidecar.
-Metadata Sidecar.
-To integrate our C2PA tooling non-intrusively with ML pipelines, the Atlas attestation client uses a sidecar container architecture.
-This allows the client to monitor the pipeline for modifications, generate manifests and update relationships through an optimized storage layer.
-In addition to collecting C2PA manifest metadata, the sidecar generates the authenticated transformation attestations for model artifacts, and tracks artifact relationships by mapping datasets to checkpoints, checkpoints to models, and maintaining configuration bindings.
-The sidecar container implementation monitors the pipeline for modifications, generating manifests that are first stored in a local cache layer before being committed to the transparency log.
-The local cache maintains an indexed hierarchy of manifests for efficient validation during pipeline execution, before the final storage in Rekor for tamper-evident provenance tracking.
-For storage optimization, our implementation decomposes manifests into constituent components.
-The assertion store, claim signatures, and metadata are stored separately, with relationships maintained through a reference system.
-This approach enables efficient updates to specific manifest components, reduced storage redundancy, optimized query performance, and scalable version tracking.
-TEE Service.
-To obtain the Intel TDX-based compute environment attestations that are included in the C2PA metadata, the sidecar interfaces with the TEE service.
-We implement a dedicated RESTful service that captures TEE state and validates runtime ML system component measurements, which are cryptographically anchored in hardware via Intel TDX.
-The TEE service implements the Confidential Containers (CoCo) Attestation format [52] for standardized claims used in TEE verification.
-We optimize attestation operations through batching and caching mechanisms.The client coordinates with the verification system through standard Kubeflow interfaces, enabling seamless integration with the other Atlas components while preserving attestation integrity.
-5.1 Kubeflow Integration
+Similarly, Supply Chain Integrity, Transparency and Trust (SCITT) [33] is an architecture for implementing distributed ledger based supply chain integrity mechanisms, providing global visibility and auditing for supply chain operations and claims.
+The SCITT architecture also includes confidential computing technologies that help ensure that only authorized parties submit claims to the transparency ledger.
+Appendix B Implementation Details
+B.1 Kubeflow Integration
 The integration with Kubeflow is achieved through custom operators and controllers that monitor pipeline execution through Kubeflow’s Metadata V2 Beta API and KFP API.
@@
 This structured approach enables verification of pipeline states while maintaining compatibility with existing workflows.
-5.2 TEE Service Claims
-For each attestation request, the client generates a JSON Web Token  [53], which adds custom claims to represent Atlas-specific information including issuer identification and public key for signature verification, temporal validity through expiration and effective timestamps, hardware TEE information and software measurements, and golden values used in verification policy enforcement
-Verification Service Implementation.
-The metadata sidecar also serves as a verification endpoint allowing pipeline components to validate artifact integrity against stored attestations.
-We optimize the performance of our staged verification system in three ways: 1) by processing changes incrementally and caching to avoid re-verifying unchanged components, 2) via batch processing of verification operations, and 3) parallel verification paths for independent component classes.
-For provenance chain validation, our implementation parallelizes verification of cross-component dependencies, version compatibility and evaluation result consistency.
+B.2 C2PA Metadata Examples
+A typical execution record captured by our system looks like:
+        "execution": {
+            "name": "training-run-132",
+            "state": "RUNNING",
+            "pipeline_spec": {
+                "parameters": {
+                    "learning_rate": 0.001,
+                    "batch_size": 32,
+                    "random_seed": 42,
+                    "optimizer_config": {
+                        "type": "Adam",
+                        "beta1": 0.9,
+                        "beta2": 0.999
+                "runtime_config": {
+                    "gcs_output_directory": "gs://...",
+                    "tensorflow_version": "2.9.0"
+For each execution, our system adds corresponding integrity measurements and verification records:
+        "integrity_measurement": {
+            "component_id": "training-run-132",
+            "tdx_quote": "base64:...",
+            "environment_hash": "sha256:...",
+            "timestamp": "2024-01-15T10:30:00Z",
+            "parameter_hash": "sha256:..."
+B.3 ML System Monitoring Procedures
+Our proof-of-concept implementation leverages several techniques to monitor ML pipeline activities with minimal intrusion into existing workflows.
+The implementation focuses on collecting runtime data about model weights, hyperparameters, and execution context while maintaining performance and compatibility with established ML frameworks.
+B.3.1 File System Monitoring in the Atlas Framework
+The Atlas sidecar implements a file system monitor written in Rust that detects checkpoint creation and modification events:
+1:procedure InitCheckpointMonitor($dir,client$)
+2:   Initialize directory and client references
+3:   Create empty checksum tracking map
+4:   Setup file system watcher
+5:end procedure
+6:procedure SetupWatcher
+7:   Create event listener for file changes
+8:   Start background monitoring thread
+9:   Register directory for change notifications
+10:end procedure
+11:procedure ScanCheckpoints
+12:   for each checkpoint file in directory do
+13:      Compute file cryptographic checksum
+14:      Store checksum in tracking map
+15:      Register existing checkpoint in metadata
+16:   end for
+17:end procedure
+18:procedure OnFileCreated($file$)
+19:   if file is checkpoint type then
+20:      Compute checksum and record creation
+21:      Update checksum tracking map
+22:   end if
+23:end procedure
+24:procedure OnFileModified($file$)
+25:   if file is checkpoint type then
+26:      Compute new checksum
+27:      Retrieve old checksum from map
+28:      if checksums differ then
+29:         Update tracking map
+30:         Record modification in metadata
+31:      end if
+32:   end if
+33:end procedure
+B.3.2 Callback/Hook Registration in PyTorch
+For our BERT Meta case study, we implemented a callback system that integrates with PyTorch’s event mechanisms. The model monitoring component operates as follows:
+1:procedure InitModelMonitor($model,client$)
+2:   Store references to model and client
+3:   Register monitoring hooks on model
+4:end procedure
+5:procedure RegisterHooks
+6:   for each layer module in model do
+7:      if module is neural network layer then
+8:         Attach forward hook for activation capture
+9:         if module has trainable weights then
+10:            Attach gradient hook for updates
+11:         end if
+12:      end if
+13:   end for
+14:end procedure
+15:procedure ForwardHook($module,input,output$)
+16:   Calculate unique layer identifier
+17:   Extract statistical metrics from output
+18:   Record activation data to metadata store
+19:end procedure
+20:procedure GradientHook($gradient$)
+21:   Calculate gradient magnitude
+22:   Record gradient event with timestamp
+23:end procedure
+Additionally,for some cases we extended PyTorch’s standard training loop with epoch-level callbacks [82]:
+1:procedure InitTrainingCallback($client$)
+2:   Store reference to metadata client
+3:end procedure
+4:procedure OnEpochStart($epoch,optimizer$)
+5:   Create hash of optimizer configuration
+6:   Record epoch start event with config hash
+7:end procedure
+8:procedure OnEpochEnd($epoch,metrics,model$)
+9:   Capture cryptographic model state snapshot
+10:   Record completion with metrics and snapshot
+11:end procedure
+These hooks operate with minimal overhead while providing comprehensive visibility into the model’s evolution.
+B.3.3 Configuration Wrappers in the Atlas Framework
+To extract hyperparameter access and modifications from PyTorch, we implement transparent wrapper classes:
+1:procedure InitTrackedConfig($config,client$)
+2:   Store config and client references
+3:   Initialize version counter to zero
+4:   Record initial configuration state
+5:end procedure
+6:procedure Get($key$)
+7:   Log access event to metadata store
+8:   Return value for requested key
+9:end procedure
+10:procedure Set($key,value$)
+11:   Retrieve current value for key
+12:   Update configuration with new value
+13:   Increment version counter
+14:   Record modification in metadata
+15:   Update configuration state hash
+16:end procedure
+17:procedure RecordState
+18:   Generate hash of current configuration
+19:   Store versioned snapshot in metadata
+20:end procedure
+21:procedure GetConfig
+22:   Return copy of configuration
+23:end procedure
+B.4 Integration Between Framework and ML Pipeline
+For the BERT Meta case study, we developed an integration layer that allows the framework components to interact with the Python ML pipeline:
+1:procedure InitBridgeService($endpoint,dir$)
+2:   Create metadata client connection
+3:   Initialize checkpoint monitor
+4:   Start bridge service
+5:end procedure
+6:procedure InitConfig($configJSON$)
+7:   Parse configuration from JSON
+8:   Create configuration tracking wrapper
+9:   Generate unique tracking identifier
+10:   return tracking identifier
+11:end procedure
+12:procedure ScanCheckpoints
+13:   Scan and register existing checkpoints
+14:end procedure
+15:procedure RecordEvent($type,data$)
+16:   if $type$ is epoch start then
+17:      Record epoch initialization
+18:   else if $type$ is epoch end then
+19:      Record epoch completion with metrics
+20:   else if $type$ is layer activation then
+21:      Record layer output statistics
+22:   else if $type$ is gradient event then
+23:      Record gradient flow information
+24:   end if
+25:end procedure
+On the Python side, we implement a complementary bridge client:
+1:procedure InitBridgeClient($socketPath$)
+2:   Connect to monitoring bridge service
+3:end procedure
+4:procedure SetupMonitoring($model,config,dir$)
+5:   Serialize configuration to JSON
+6:   Initialize configuration tracking
+7:   Scan existing model checkpoints
+8:   Create model monitoring hooks
+9:   Setup training loop callbacks
+10:   return monitoring components
+11:end procedure
+12:procedure RecordLayerData($layer,stats$)
+13:   Package activation statistics
+14:   Send to bridge as layer event
+15:end procedure
+16:procedure RecordGradient($magnitude,time$)
+17:   Package gradient information
+18:   Send to bridge as gradient event
+19:end procedure
+20:procedure RecordEpochStart($epoch,optHash$)
+21:   Package epoch initialization data
+22:   Send to bridge as epoch start event
+23:end procedure
+24:procedure RecordEpochEnd($epoch,metrics,hash$)
+25:   Package completion metrics
+26:   Send to bridge as epoch end event
+27:end procedure
+Figure 2: Altas sidecar collector showing data flow between Python ML environment and framework: The diagram illustrates how monitoring events are cached before being committed to the transparency log.
+This hybrid architecture enables monitoring of the BERT training process with minimal modifications to the existing pipeline, while leveraging efficient system-level operations for the monitoring infrastructure.
+B.5 Attestation Client Storage Optimization
+As a storage optimization, the attestation client’s metadata sidecar first stores all generated C2PA manifests in a local cache layer before being committed to the transparency log.
+The local cache maintains an indexed hierarchy of manifests for efficient validation during pipeline execution, before final storage in Rekor for tamper-evident provenance tracking.
+More specifically, we decompose manifests into constituent components within the cache.
+The C2PA metadata assertions, claim signatures, and pipeline metadata are stored separately, with relationships maintained through a reference system.
+This approach enables efficient updates to specific manifest components, reduced storage redundancy, optimized query performance, and scalable version tracking.
+B.6 Verification Service Optimizations
+Figure 3: Verification workflow implementation for Atlas components.
+The system classifies artifacts and processes changes incrementally, preserving cached states for unchanged components.
+Related artifacts are grouped for batch processing, with parallel validation of relationships.
+During errors, only affected components are invalidated, reducing verification overhead while maintaining security guarantees.
+For provenance chain validation, our verification service implementation parallelizes verification of cross-component dependencies, version compatibility and evaluation result consistency.
 The system also maintains verification checkpoints that serve as trusted reference points, enabling partial verification from the last known good state instead of complete chain recomputation.
@@
 Preliminary testing shows these optimizations reduce verification time by up to 50% through parallelization and caching, while maintaining security guarantees.
-5.3 Verification Flow
-Figure 2: Verification workflow implementation showing how components are classified and processed: Changes trigger incremental validation that preserves cached states for unchanged components.
-The system groups related artifacts (training data, model weights, configurations) for batch processing, validates their relationships in parallel, and selectively invalidates only affected components on failures.
-This approach reduces verification overhead while maintaining security guarantees.
-5.4 Framework Adaptability
+B.7 Framework Adaptability
 The BERT deployment validated our framework’s flexibility across different ML environments.
@@
 Custom adapters enabled integration without modifying existing ML infrastructure, demonstrating the framework’s ability to enhance security while preserving established workflows.
-6 Evaluation
-We validate our framework through a security analysis, preliminary performance testing and two case studies: BERT Meta [3, 4] and BGE reranker [5].
-6.1 Security Analysis
-Atlas provides measures against the threats outlined in §3.3 through multiple security mechanisms.
-For MLaaS provider threats, the hardware-rooted TEEs in Atlas isolate sensitive computations and detect malicious insider tampering with executing ML pipelines.
-The attestation client continuously validates the runtime environment, generating ML system measurements that are cryptographically bound to model artifacts.
-Against hub threats, the Atlas verification system ensures artifact integrity through cryptographic measurements and signatures.
-The provenance chain maintains verifiable links between artifacts, preventing mismatched signatures or compromised dependencies from propagating.
-To address artifact producer threats, Atlas relies on comprehensive provenance tracking that captures all dependencies and modifications.
-The transparency layer provides an immutable record of pipeline operations, enabling detection of undeclared dependencies or intentional omissions.
-6.2 Preliminary Performance Analysis
-Our implementation addresses the computational overhead of security mechanisms through several optimization strategies.
-Preliminary tests demonstrate that Atlas’ security mechanisms introduce minimal overhead, with training overhead remaining under 8% across all test cases.
-The verification processes scale linearly with model size, while our caching strategies (see §5) effectively optimize performance for iterative ML workflows; we observe a reduction of up to 50% of verification latency compared to the unoptimized implementation, and near-constant time for cached component verification.
-For large-scale operations, parallel processing capabilities maintain performance at scale through: 1) concurrent verification operations when dependencies allow; 2) cache optimization to reduce repeated verification costs; 3) selective cache invalidation for error handling.
-We plan to conduct more extensive performance benchmarking as part of future work.
-6.3 Case Studies
-BERT was selected for its complex architecture and widespread production use.
-Our implementation covers the complete lifecycle from pre-trained model fine-tuning through deployment, with isolated execution environments for attention computations and weight updates.
-To validate Atlas’ adaptability, we also implement security controls for BGE reranker on Intel Xeon processors.
-Atlas secures both instruction-based and layerwise reranking configurations, with training data structured as JSON records containing query-positive-negative text triplets.
-The provenance system tracks model adaptations across both base model variations (Gemma-2B and MiniCPM-2B), maintaining verifiable records of hyperparameters and training progressions.
-Both implementations demonstrate performance characteristics consistent with our analysis in  §6.2, while preserving training efficiency and security guarantees.
-7 Case Study: BERT Pipeline
-Our proof-of-concept implementation demonstrates the framework’s capabilities through securing a BERT model development and deployment pipeline.
-We chose BERT Meta [3, 4] for this case study due to its complex architecture, widespread use in production environments, and frequent fine-tuning requirements that test our framework’s adaptability.
-The implementation covers the complete model lifecycle from pre-trained model adaptation through deployment.
-For BERT’s transformer architecture, we implemented isolated execution environments for attention computations and weight updates.
-The hardware-backed attestation system provides continuous runtime verification during fine-tuning operations, while our provenance system maintains verifiable records of all model evolution steps from the base model through task-specific adaptations.
-Performance analysis shows practical viability with training overhead under 8%, primarily from manifest generation, and verification processes scaling linearly with model size.
-The system’s parallel processing capabilities and cache optimization strategies maintained performance at scale.
-8 Discussion and Conclusion
-Our implementation and evaluation reveal several key considerations for practical deployment ML pipeline security.
-Organizations must balance security requirements with operational efficiency, considering factors like verification frequency, attestation depth, and computational overhead.
-Atlas’ modular design enables customization of these tradeoffs while maintaining core security guarantees.
-Current hardware security architectures present integration challenges, particularly around GPU and accelerator attestation support.
-While Atlas provides secure boundaries for CPU-GPU data transfers, emerging technologies like confidential GPU computing will enable more comprehensive coverage.
-Integration with legacy HPC systems also requires careful consideration, as traditional job schedulers often lack modern observability capabilities needed for fine-grained monitoring and attestation.
-9 Adoption Considerations
-9.1 Storage Optimizations
-9.1.1 Storage and Scalability
+Appendix C Adoption Considerations
+C.1 Storage Optimizations
+C.1.1 Storage and Scalability
 Our analysis suggests opportunities for optimizing manifest storage through decomposed and hybrid architectures.
@@
 Organizations should consider distributed storage strategies that balance immutability requirements with query performance needs.
-9.1.2 Decomposed Storage Model
+C.1.2 Decomposed Storage Model
 Instead of storing complete manifests as single documents, the system could decompose manifests into their constituent components.
@@
   • Link components through cryptographic references preserved in the ledger
-9.1.3 Hybrid Storage Architecture
+C.1.3 Hybrid Storage Architecture
 A hybrid approach could maintain critical verification data in Rekor for its transparency guarantees while storing detailed manifest data in optimized storage systems.
@@
 Performance testing indicates potential reduction in storage requirements and query latency through these alternative approaches.
-9.2 Deployment Guidelines
-9.2.1 Organizational Adaptations
+C.2 Deployment Guidelines
+C.2.1 Organizational Adaptations
 Organization-specific adaptations are necessary to align with existing infrastructure and security policies. Key considerations include:
@@
   • Compliance with organizational security standards
-9.2.2 Security Requirement Balance
+C.2.2 Security Requirement Balance
 Security requirement balance directly impacts operational efficiency.
@@
 A more balanced approach might implement full verification at critical pipeline stages while using lightweight checks during intermediate steps.
-9.2.3 Computational Overhead Management
+C.2.3 Computational Overhead Management
 Computational overhead management becomes crucial when scaling the framework across large ML operations.
@@
   • Parallel verification processing where possible - This approach utilizes available computational resources effectively by running verification operations concurrently when component dependencies allow.
-10 Future Work
+Appendix D Future Work
 Several potential enhancements could extend our framework’s capabilities and applicability:
@@
 These properties would enhance the framework’s ability to detect subtle manipulations that might not affect model hashes but could impact model behavior.
-Our work demonstrates that robust security measures can be integrated into ML lifecycles without significant performance impact.
-The combination of hardware-backed security guarantees with provenance tracking provides a foundation for securing ML pipelines.
-Our case studies show Atlas’ ability to maintain strong security properties while preserving established development practices.
-We identify three directions for further enhancing ML pipeline security: 1.
-extending hardware security coverage to specialized ML accelerators will enable more comprehensive protection of ML computations; 2.
-standardizing provenance tracking across the industry will significantly improve interoperability and adoption; 3.
-developing robust algorithmic verification methods and model guardrails will strengthen protection against sophisticated attacks targeting model behavior, allowing for flexible integration of emerging defenses.
+Throughout this process, Atlas provides tamper-evident records through its transparency log and TEE-based attestations.
+The verification service requires access to both the model artifacts and the cryptographic measurements in the transparency log to confirm the integrity of the complete ML pipeline, ensuring that no unauthorized modifications occurred during the model’s lifecycle.